### v1

# supplement

Abstract
This paper presents several forecasting methods to model and forecast subnational age distribution of death counts. The age distribution of death counts has many similarities to probability density functions, which are nonnegative and have a constrained integral, and thus live in a constrained nonlinear space. To address the nonlinear nature of objects, we implement a cumulative distribution function transformation that has an additional monotonicity. Using the Japanese subnational life-table death counts obtained from the Japanese Mortality Database (2025), we evaluate the forecast accuracy of the transformation and forecasting methods. The improved forecast accuracy of life-table death counts implemented here will be of great interest to demographers in estimating regional age-specific survival probabilities and life expectancy.
Keywords: constrained functional time series; cumulative distribution function transformation; (split) conformal prediction; life-table death counts; principal component analysis
1 Introduction
Density-valued objects for multiple groups are common, with examples including income distributions across different populations (Kneip & Utikal, 2001), financial return distributions for multiple stocks (Petersen et al., 2022), distributions of bidding times in online auctions for various items (Wang et al., 2008), and subnational age distributions of deaths in demography (Jiménez-Varón et al., 2025), among others. We present several forecasting methods to model and predict the subnational age distribution of deaths, which are also useful for applications in actuarial science, such as fixed-term or lifetime annuity pricing.
Since age distributions of deaths naturally take the form of density functions, they are well-suited for modeling using advanced forecasting techniques from compositional and functional data analyses. Oeppen (2008) demonstrates that using compositional data analysis to forecast the age distribution of death counts performs similarly in accuracy to forecasting age-specific mortality rates. Within the compositional data analysis, the Lee-Carter model is frequently applied to model and forecast unconstrained data. Shang & Haberman (2020) extend this approach by incorporating a functional data analytic perspective, which includes multiple principal components and nonparametric smoothing. Recently, Shang & Haberman (2025) introduce a cumulative distribution function (CDF) transformation, converting each year’s age distribution of deaths into a probability bounded within the unit interval to address the presence of zero values. Through cumulative summation, a probability density function can be transformed into a CDF, which provides the additional benefit of monotonicity (see also Mayhew & Smith, 2013). With a time series of CDFs, it is common to model and extrapolate its pattern via a logistic transformation.
Multiple density-valued objects are observed over time are related to high-dimensional functional time series (HDFTS). In the statistical literature, Zhou & Dette (2023) derive Gaussian and multiplier bootstrap approximations for sums of HDFTS. With these approximations, they construct joint simultaneous confidence bands for the mean functions and develop a hypothesis to test whether the mean functions in the cross-sectional dimension exhibit parallel behavior. Hallin et al. (2023) investigate the representation of HDFTS using a factor model, determining conditions on the eigenvalues of the covariance operator crucial for establishing the existence and uniqueness of the factor model. Gao et al. (2019) adopt a two-stage approach, combining truncated principal component analysis and a separate scalar factor model for the resulting panels of scores. Tavakoli et al. (2023) introduce a functional factor model with a functional factor loading and a vector of real-valued factors, while Guo et al. (2024) propose a functional factor model with a real-valued factor loading and a functional factor. Leng et al. (2024) introduce a unified framework that accommodates both types of factor models. Tang et al. (2022) study clustering for age-specific subnational mortality rates, which is an example of HDFTS. Li et al. (2024) introduce hypothesis tests and estimation procedures for testing and locating (common) change points. Chang et al. (2025) develop a two-stage procedure for modeling and forecasting HDFTS.
Our contributions are twofold. First, we present several visualization techniques to capture the patterns in age distribution of death counts and display contrasts between subnational and national data. Second, we revisit several forecasting methods and compare their point and interval forecast accuracy with applications to subnational life-table death counts.
Our paper is structured as follows. In Section 2, we describe our motivating data set and introduce a series of image plots for displaying important features in HDFTS. In Section 3, we present the CDF transformations. Within the CDF transformation, we consider a number of forecasting methods in Section 4 to model and forecast unconstrained functional time-series data. In Section 5, we propose two general strategies for constructing pointwise prediction intervals for the age distribution of death counts. In Section 6, we evaluate and compare point forecast accuracy using the Kullback-Leibler divergence (KLD) and Jensen-Shannon divergence (JSD), and interval forecast accuracy using the coverage probability difference (CPD) and interval score. Section 7 concludes with some ideas on how the methodology presented can be further extended.
2 Subnational age distribution of death in Japan
In many developed countries such as Japan, increases in longevity and an aging population have led to concerns about the sustainability of pensions, health, and aged care systems (see, e.g., Coulmas, 2007). These concerns have resulted in a surge of interest among government policy makers in accurately modeling and forecasting age-specific mortality. Subnational forecasts of age-specific mortality are useful for informing policy within local regions and any improvement in forecast accuracy of mortality is beneficial for determining the allocation of current and future resources at the national and subnational levels.
In demography, the age distribution of deaths provides important insights into longevity and lifespan variability that cannot be grasped directly from the central mortality rate or survival function. As pointed out by Oeppen (2008), the age distribution of death counts is more suitable than central mortality rates for computing life expectancy and annuity premia. In addition to providing an informative description of the mortality experience of a population, life-table death counts yield readily available information on the “central longevity indicators” (e.g., mean, median, and modal age at death, see Cheung et al., 2005; Canudas-Romo, 2010) and lifespan variability (e.g., van Raalte & Caswell, 2013; Aburto & van Raalte, 2018).
We study Japanese lifetable death counts from 1975 to 2022, obtained from the Japanese Mortality Database (2025). We consider ages from 0 to 110 in single years of age. Figure 1 shows rainbow plots of the female and male age-specific life-table death counts in Japan from 1975 to 2022. The time ordering of the curves follows the color order of a rainbow, where curves from the distant past are shown in red and the more recent curves are shown in violet. The figures show typical mortality curves for a developed country, with a decreasing trend in infant death counts. A typical negatively skewed distribution for the life-table death counts is apparent where the peaks shift to higher ages for both females and males. This shift is a main source of longevity risk, which is a major issue for insurers and pension funds, especially in the selling and risk management of annuity products (see Denuit et al., 2007, for a discussion). Moreover, the spread of the distribution indicates lifespan variability. A decrease in variability over time can be observed directly and can be measured by the Gini coefficient (Wilmoth & Horiuchi, 1999; Debón et al., 2017).
(a) Japanese female data
(b) Japanese male data
Figure 1: Functional time series graphical displays of age-specific life-table death counts from 1975 to 2022 in a single-year group. The oldest years are shown in red, with the most recent years in violet. Curves are ordered chronologically according to the colors of the rainbow.
2.1 Image plots
Another visual perspective of the data is shown in the image plot of Figure 2. Here, we graph the symmetric KLD between life-table death counts for each prefecture to life-table death counts for the whole country, thus allowing relative mortality comparisons to be made. A divergent color palette is used with red representing larger KLD values (i.e., greater difference between subnational and national data) and yellow denoting smaller KLD values. The prefectures are ordered geographically from north to south, so the most northerly prefecture (Hokkaido) is at the top of the panels, and the most southerly prefecture (Okinawa) is at the bottom.
Figure 2: Image plots showing the symmetric KLD introduced in (9) between life-table death counts of a prefecture and life-table death counts of the whole country. The top panel shows mortality averaged over ages, while the bottom panel shows mortality averaged over years. Prefectures are numbered geographically from north to south.
The top row of the panels shows mortality for each prefecture and year, averaged over all ages. In 2011, in Iwate, Miyagi, and Fukushima, there was a large increase in mortality compared to other prefectures. These are northern coastal regions, and the inflated relative mortality is due to the tsunami of 11 March 2011 (see, e.g., Nakahara & Ichikawa, 2013) caused by a magnitude–9.0 earthquake.
In 1995, there was an increase in mortality for Hyogo, which corresponds to the Kobe (Great Hanshin) earthquake (magnitude 6.9) of 17 January 1995. Also evident is the mortality difference in Okinawa up to 2000, suggesting a decline in the relative mortality advantages enjoyed by residents in this region.
The bottom of the panels shows mortality for each prefecture and age, averaged over all years. Several striking features are apparent. There are strong differences between the prefectures for children and teenagers; this is possibly due to socio-economic differences and the accessibility of health services.
2.2 Functional cross-correlation plot
Let $d_{t,s}^{g}(u)$ be the age distribution of deaths for age $u$, prefecture $s$, gender $g$ at year $t$, respectively. There are 111 ages, $S=47$ prefectures, and $48$ years. Following Kokoszka et al. (2017) and Mestre et al. (2021), the functional autocorrelation coefficient at lag $h$ can be defined as
[…]
where $h$ denotes a lag and $C_{h}(u,u)=\text{cov}[d_{t,s}^{g}(u),d_{t+h,s}^{g}(u)]$. The functional cross-covariance at lag $h$ can be expressed as
[…]
where $C_{h}[d_{s}^{g}(u),d_{\text{national}}^{g}(v)]=\text{cov}[d_{t,s}^{g}(u),d_{t+h,\text{national}}^{g}(v)]$. The functional cross-correlation at lag $h$ is then given by
[…]
When $d^{g}_{s}(u)=d_{\text{national}}^{g}(v)$, cross correlation in (2) reduces to autocorrelation in (1).
In Figure 3, we display the functional cross-correlation functions showing the dependence of life-table death counts between each prefecture and the whole country. Each subnational data has a comparably stronger cross correlation with the national data at lags 1 and 2. The cross-correlation reduces as lag increases.
Figure 3: Functional cross-correlation plots showing the dependence of the female, male, and total life-table death counts between each prefecture and the whole country. Prefectures are numbered geographically from north to south.
3 Cumulative distribution function transformation
In the CDF transformation, we first convert an age distribution of death for each year to a probability by dividing its radix $10^{5}$. Via the cumulative sum, we transform a probability density function into a CDF,
$D_{t,s}^{g}(u_{x})=\sum^{x}_{i=1}d_{t,s}^{g}(u_{i}),\quad x=1,\dots,111,\quad t=1,\dots,n,$
where $D_{t,s}^{g}(u_{111})=1$. In doing so, it enjoys an additional benefit of monotonicity (see also Mayhew & Smith, 2013).
Then, we perform a logistic transformation,
$\mathcal{X}_{t,s}^{g}(u_{x})=\text{logit}[D_{t,s}^{g}(u_{x})]=\ln\Big{[}\frac{D_{t,s}^{g}(u_{x})}{1-D_{t,s}^{g}(u_{x})}\Big{]},$
where $\ln(\cdot)$ denotes the natural logarithm. Since $D_{t,s}^{g}(u_{111})=1$, the last column is removed to avoid the undefinedness of the logistic transformation. By considering age as a continuum, the object $\mathcal{X}_{t,s}^{g}(u)$ is an example of a high-dimensional functional time series.
To transform $\mathcal{X}_{t,s}^{g}(u_{x})$ back to the original scale, we first perform an inverse logit transformation, obtaining
$D_{t,s}^{g}(u_{x})=\frac{\exp^{\mathcal{X}_{t,s}^{g}(u_{x})}}{1+\exp^{\mathcal{X}_{t,s}^{g}(u_{x})}},$
where we add back the last column of ones. Then, we take the first-order differencing to obtain
$d_{t,s}^{g}(u_{x})=\Delta_{i=1}^{x}D_{t,s}^{g}(u_{i})\times 10^{5},$
where $\Delta$ represents the first-order differencing, where $\Delta_{i=1}^{1}D_{t,s}^{g}(u_{1})=D_{t,s}^{g}(u_{1})$. The constant $10^{5}$ is the radix of the life-table death count. Since the data are observed discretely, we use the cumulative sum and first-order differencing as approximations for integration and differentiation, respectively.
4 Functional time-series forecasting models
[…]
4.1 Univariate functional time-series (UFTS) forecasting method
We assume $\mathcal{X}_{t,s}^{g}(u)$ has a finite mean and variance. Then, its non-negative definite covariance function is given by
$c_{s}^{g}(u,v)$ $:=\text{Cov}[\mathcal{X}_{s}^{g}(u),\mathcal{X}_{s}^{g}(v)]$
$=\text{E}\left\{[\mathcal{X}_{s}^{g}(u)-\mu_{s}^{g}(u)][\mathcal{X}_{s}^{g}(v)-\mu_{s}^{g}(v)]\right\},\qquad\forall u,v\in\mathcal{I},$
where $\mu_{s}^{g}(u)=\text{E}\{\mathcal{X}_{t,s}^{g}(u)\}$. Via Mercer’s lemma, there exists a set of orthonormal sequence $(\phi_{k})$ of continuous functions in $\mathcal{L}^{2}(\mathcal{I})$ and a non-increasing sequence $\lambda_{k,s}^{g}$ of positive numbers, such that
$c_{s}^{g}(u,v)=\sum^{\infty}_{k=1}\lambda_{k,s}^{g}\phi_{k,s}^{g}(u)\phi_{k,s}^{g}(v),$
where $\lambda_{k,s}^{g}$ and $\phi_{k,s}^{g}(u)$ denote the $k$th eigenvalue and eigenfunction, respectively. The Karhunen-Loève expansion of a stochastic process $\mathcal{X}_{t,s}^{g}(u)$ can be expressed as
$\mathcal{X}_{t,s}^{g}(u)$ $=\mu_{s}^{g}(u)+\sum_{k=1}^{\infty}\beta_{t,k,s}^{g}\phi_{k,s}^{g}(u)$
$=\mu_{s}^{g}(u)+\sum_{k=1}^{K}\beta_{t,k,s}^{g}\phi_{k,s}^{g}(u)+e_{t,s}^{g}(u),$
where the principal component score $\beta_{t,k,s}^{g}$ is given by the projection of $[\mathcal{X}_{t,s}^{g}(u)-\mu_{s}^{g}(u)]$ in the direction of the $k$th eigenfunction $\phi_{k,s}^{g}(u)$, i.e., $\beta_{t,k,s}^{g}=\langle\mathcal{X}_{t,s}^{g}(u)-\mu_{s}^{g}(u),\phi_{k,s}^{g}(u)\rangle$; and $e_{t,s}^{g}(u)$ denotes the model error function with a mean of zero and a finite variance, and $K<n$ is the number of retained components.
There are a number of ways to select the retained number of principal components, such as the bootstrap approach proposed by Hall & Vial (2006) and Bathia et al. (2010), description length approach proposed by Poskitt & Sengarapillai (2013), pseudo-Akaike information criterion (Shibata, 1981), scree plot (Cattell, 1966), and eigenvector variability plot (Tu et al., 2009). We consider the eigenvalue ratio criterion (EVR) of Li et al. (2020), defined as
$K=\operatorname*{arg\,min}_{k=1,\dots,k_{\max}}\left\{\frac{\widehat{\lambda}_{k+1,s}^{g}}{\widehat{\lambda}_{k,s}^{g}}\times\mathds{1}\Big{\{}\frac{\widehat{\lambda}_{k,s}^{g}}{\widehat{\lambda}_{1,s}^{g}}\geq\eta\Big{\}}+\mathds{1}\Big{\{}\frac{\widehat{\lambda}_{k,s}^{g}}{\widehat{\lambda}_{1,s}^{g}}<\eta\Big{\}}\right\},$ (3)
where $\mathds{1}\{\cdot\}$ denotes the binary indicator function, $\eta=\frac{1}{\ln[\max(\widehat{\lambda}_{1,s}^{g},n)]}$ is a small positive value. Instead of searching through $n$, we restrict our searching range by setting $k_{\max}=\#\{k|\widehat{\lambda}_{k,s}^{g}\geq\frac{1}{n}\sum^{n}_{k=1}\widehat{\lambda}_{k,s}^{g},k\geq 1\}$. For comparison, we also consider $K=6$ as suggested in Hyndman et al. (2013).
Conditioning on the observed functions $\bm{\mathcal{X}}_{s}^{g}(u)=[\mathcal{X}_{1,s}^{g}(u),\dots,\mathcal{X}_{n,s}^{g}(u)]$ and the estimated functional principal components $\bm{B}=\big{\{}\phi_{1,s}^{g}(u),\dots,\phi_{K,s}^{g}(u)\big{\}}$, the $h$-step-ahead point forecast of $\mathcal{X}_{n+h,s}^{g}(u)$ can be expressed as
$\widehat{\mathcal{X}}_{n+h|n,s}^{g}(u)$ $=\text{E}\big{[}\mathcal{X}_{n+h,s}^{g}(u)|\bm{\mathcal{X}}_{s}^{g}(u),\bm{B}\big{]}$
$=\widehat{\mu}_{s}^{g}(u)+\sum^{K}_{k=1}\widehat{\beta}_{n+h|n,k,s}^{g}\widehat{\phi}_{k,s}^{g}(u),$
where $\widehat{\mu}_{s}^{g}(u)=\frac{1}{n}\sum^{n}_{t=1}\mathcal{X}_{t,s}^{g}(u)$ and $\widehat{\beta}_{n+h|n,k,s}^{g}$ denotes time-series forecasts of the $k$th principal component scores. The forecasts of these scores can be obtained via a univariate time-series forecasting method, such as exponential smoothing (Hyndman et al., 2008). Computationally, we implement the ets function in the forecast package (Hyndman & Khandakar, 2008), which selects the optimal parameters based on the corrected Akaike information criterion.
4.2 Multivariate functional time-series (MFTS) forecasting method
The univariate functional time-series method does not incorporate a correlation between female and male data within the same region. However, explicitly modeling the correlation between multiple series may improve forecast accuracy. We consider a multivariate functional time-series method to jointly model and forecast multiple series that could be correlated with each other.
Let $\mathcal{X}_{t,s}^{F}(u)$ and $\mathcal{X}_{t,s}^{M}(u)$ represent female and male subnational age distribution of death counts. As our multiple functional time series have the same function support, we consider data where each observation consists of $\omega\geq 2$ functions, that is, $[\mathcal{X}_{t,s}^{F}(u),\mathcal{X}_{t,s}^{M}(u)]\in\mathds{R}^{2}$, where $u\in\mathcal{I}$.
The multivariate functional time series are stacked in a vector. Let $\nu^{F}_{s}(u)$ and $\nu^{M}_{s}(u)$ represent the mean function. For $u,v\in\mathcal{I}$, the theoretical cross-covariance function can be defined with elements
$c_{s}(u,v):=\text{Cov}[\mathcal{X}^{F}_{s}(u),\mathcal{X}^{M}_{s}(v)]=\text{E}\left\{[\mathcal{X}^{F}_{s}(u)-\nu^{F}_{s}(u)][\mathcal{X}^{M}_{s}(u)-\nu^{M}_{s}(u)]\right\}.$
The Karhunen-Loève expansion of a stochastic process can be expressed as
$\bm{\mathcal{X}}_{s}(u)=\bm{\nu}_{s}(u)+\bm{\Phi}_{s}(u)\bm{\beta}_{t,s}^{\top},$
where $\bm{\mathcal{X}}_{s}(u)=[\bm{\mathcal{X}}_{s}^{F}(u),\bm{\mathcal{X}}_{s}^{M}(u)]$, $\bm{\mathcal{X}}_{s}^{F}(u)=[\mathcal{X}_{1,s}^{F}(u),\dots,\mathcal{X}_{n,s}^{F}(u)]$ denote stacked historical functions, and $\bm{\nu}_{s}(u)=[\bm{\nu}_{s}^{F}(u),\bm{\nu}_{s}^{M}(u)]$. Moreover, $\bm{\beta}_{s}=(\bm{\beta}_{1,s},\dots,\bm{\beta}_{K,s})$ and $\bm{\beta}_{1,s}=(\beta_{1,1,s},\dots,\beta_{n,1,s})$ is the vector of principal component scores, and
$\bm{\Phi}_{s}(u)=\left[\begin{array}[]{cccccc}\phi_{1,s}^{F}(u)&\cdots&\phi_{K,s}^{F}(u)&&\cdots&\\ &\cdots&&\phi_{1,s}^{M}(u)&\cdots&\phi_{K,s}^{M}(u)\end{array}\right]_{2\times(K\times 2)}$
where $K$ denotes the retained number of principal components shared by females and males.
Conditioning on the past functions $\bm{\mathcal{X}}_{s}(u)$ and the estimated functional principal components $\bm{\Phi}_{s}(u)$, the $h$-step-ahead point forecast of $\bm{\mathcal{X}}_{n+h,s}(u)$ can be expressed as
$\widehat{\bm{\mathcal{X}}}_{n+h|n,s}(u)=\text{E}\left[\bm{\mathcal{X}}_{n+h,s}(u)|\bm{\mathcal{X}}_{s}(u),\bm{\Phi}_{s}(u)\right]=\widehat{\nu}_{s}(u)+\widehat{\bm{\Phi}}_{s}(u)\widehat{\bm{\beta}}_{n+h|n,s}^{\top},$
where $\widehat{\bm{\beta}}_{n+h|n,s}$ denotes the time-series forecasts of the principal component scores corresponding to the female and male series, $\widehat{\nu}_{s}(u)$ and $\widehat{\bm{\Phi}}_{s}(u)$ denote the estimated mean function and estimated functional principal components, respectively.
4.3 Multilevel functional time-series (MLFTS) forecasting method
The multilevel functional data model strongly resembles to the two-way functional analysis of variance studied by Morris & Carroll (2006) and Cuesta-Albertos & Febrero-Bande (2010). The basic idea is to extract common patterns shared by multiple series $R_{t,s}(u)$ and series-specific patterns $U_{t,s}^{g}(u)$. The common trend and series-specific trend are modeled by projecting them onto the eigenvectors of covariance functions of the aggregated and series-specific stochastic processes, relatively. For $t=1,\dots,n$, a curve can be expressed as
$\mathcal{X}_{t,s}^{g}(u)=\mu^{g}_{s}(u)+R^{c}_{t,s}(u)+U_{t,s}^{g}(u),$ (4)
where $R^{c}_{t,s}(u)$ can be the simple average of the female and male series. To ensure model and parameter identifiability, we implement a two-stage estimation procedure based on functional principal component decomposition.
Since the stochastic processes $R^{c}_{t,s}(u)$ and $U^{g}_{t,s}(u)$ are unknown in practice, the population eigenvalues and eigenfunctions can only be approximated at best through a set of realizations $\bm{R}^{c}_{s}(u)$ and $\bm{U}_{s}^{g}(u)$. From the covariance function of $\bm{R}^{c}_{s}(u)$, we can extract a set of functional principal components and their associated scores, along with a set of residual functions. From the covariance function of the residuals, we can then extract a second set of functional principal components and their associated scores. While the first functional principal component decomposition captures the common pattern across multiple series, the second functional principal component decomposition captures the series-specific residual trend.
The sample versions of series-specific mean function, a common trend and series-specific residual trend for a set of functional time series can be estimated by
$\widehat{\mu}_{s}^{g}(u)$ $=\frac{1}{n}\sum^{n}_{t=1}\mathcal{X}_{t,s}^{g}(u)$ (5)
$R_{t,s}^{c}(u)$ $\approx\sum^{K}_{k=1}\beta_{t,k,s}^{c}\phi_{k,s}^{c}(u)$ (6)
$U_{t,s}^{g}(u)$ $\approx\sum^{L}_{l=1}\gamma_{t,l,s}^{g}\psi_{l,s}^{g}(u),$ (7)
where $\widehat{\mu}^{g}_{s}(u)$ represents the simple average of the female or male series, $\bm{\beta}_{k,s}^{c}=(\beta_{1,k,s}^{c},\dots,\beta_{n,k,s}^{c})$ represents the $k$th sample principal component scores of $\bm{R}^{c}_{s}(u)$, and $\bm{\gamma}_{l,s}^{g}=(\gamma_{1,l,s}^{g},\dots,\gamma_{n,l,s}^{g})$ represents the $l$th sample principal component scores of $\bm{U}_{s}^{g}(u)$.
Plugging (5) to (7) into (4), we obtain
$\mathcal{X}_{t,s}^{g}(u)\approx\widehat{\mu}^{g}_{s}(u)+\sum^{K}_{k=1}\beta_{t,k,s}^{c}\phi_{k,s}^{c}(u)+\sum^{L}_{l=1}\gamma_{t,l,s}^{g}\psi_{l,s}^{g}(u)+e_{t,s}^{g}(u),$
where $e_{t,s}^{g}(u)$ denotes measurement error with a finite variance.
To select the retained number of components, we use the eigenvalue ratio criterion in (3). A feature of the multilevel method is its ability to estimate the proportion of variability explained by aggregated data. For a given population $g$, a measure of the within-cluster variability is given by
$\frac{\sum_{k=1}^{K}\widehat{\lambda}_{k,s}^{c}}{\sum_{k=1}^{K}\widehat{\lambda}_{k,s}^{c}+\sum_{l=1}^{L}\widehat{\lambda}_{l,s}^{g}}.$
When the common trend can explain the primary mode of total variability, the value of the within-cluster variability is close to 1.
[…]
$\widehat{\mathcal{X}}_{n+h|n,s}^{g}(u)$ $=\text{E}[\mathcal{X}_{n+h,s}^{g}(u)|\mu^{g}_{s}(u),\bm{\mathcal{X}}^{g}_{s}(u),\bm{\Phi}_{s}(u),\bm{\Psi}_{s}^{g}(u)]$
$=\widehat{\mu}^{g}_{s}(u)+\sum^{K}_{k=1}\widehat{\beta}_{n+h|n,k,s}^{c}\widehat{\phi}_{k,s}^{c}(u)+\sum^{L}_{l=1}\widehat{\gamma}_{n+h|n,l,s}^{g}\widehat{\psi}_{l,s}^{g}(u),$
where $\widehat{\mu}_{s}^{g}(u)=\frac{1}{n}\sum^{n}_{t=1}\mathcal{X}_{t,s}^{g}(u)$, $\{\widehat{\beta}_{n+h|n,k,s}^{c},k=1,\dots,K\}$ and $\{\widehat{\gamma}_{n+h|n,l,s}^{g},l=1,\dots,L\}$ are the forecasted principal component scores, obtained from a univariate time-series forecasting method.
4.4 Two-way functional analysis of variance
To model $\mathcal{X}_{t,s}^{g}(u)$, we consider a two-way functional ANOVA decomposition, expressed as:
$\mathcal{X}_{t,s}^{g}(u)=\mu(u)+\alpha_{s}(u)+\beta^{g}(u)+\epsilon_{t,s}^{g}(u),$
where $\mu(u)$ denotes the functional grand effect, $\alpha_{s}(u)$ denotes the functional row effect, $\beta^{g}(u)$ denotes the functional column effect, and $\epsilon_{t,s}^{g}(u)$ denotes the error function.
The functional ANOVA model can be estimated by means with
$\widehat{\mu}(u)$ $=\frac{1}{n_{s}\times 2\times n}\sum^{n_{s}}_{s=1}\sum^{2}_{g=1}\sum^{n}_{t=1}\mathcal{X}_{t,s}^{g}(u)$
$\widehat{\alpha}_{s}(u)$ $=\frac{1}{2\times n}\sum^{2}_{g=1}\sum^{n}_{t=1}\mathcal{X}_{t,s}^{g}(u)-\widehat{\mu}(u)$
$\widehat{\beta}^{g}(u)$ $=\frac{1}{n_{s}\times n}\sum^{n_{s}}_{s=1}\sum^{n}_{t=1}\mathcal{X}_{t,s}^{g}(u)-\widehat{\mu}(u).$
Some identifiability constraints exist, $\forall u\in\mathcal{I}$, $\sum^{n_{s}}_{s=1}\alpha_{s}(u)=\sum^{2}_{g=1}\beta^{g}(u)=0$, and $\sum^{n_{s}}_{s=1}\mathcal{X}_{t,s}^{g}(u)=\sum^{2}_{g=1}\mathcal{X}_{t,s}^{g}(u)=0$, $\forall t$.
The residual functions $\widehat{\epsilon}_{t,s}^{g}(u)=\mathcal{X}_{t,s}^{g}(u)-\widehat{\mu}(u)-\widehat{\alpha}_{s}(u)-\widehat{\beta}^{g}(u)$ are time-varying, for $t=1,\dots,n$. Following Jiménez-Varón et al. (2024), we model and forecast the residual functions via the MFTS.
4.5 Two-stage functional principal component analyses
Gao et al. (2019) propose a two-stage functional principal component decomposition to model and forecast HDFTS. One of the earliest methods for comparison (see, e.g., Tavakoli et al., 2023; Shang, 2025), it consists of three parts.
  1. 1)
    Functional principal component analysis is performed separately on each set of functional time series, resulting in $P$ sets of principal component scores of low dimension $p_{0}$.
  2. 2)
    The first functional principal component scores from each of $P$ sets of functional time series are combined into an $P\times 1$ vector. We fit factor models to the scores to further reduce the dimension into an $r\times 1$ vector, where $r<<P$. The same is carried out for the second, third, and so on until the $p_{0}$th scores. The vector of $P$ functional time series is reduced to an $r\times p_{0}$ matrix consisting of factors.
  3. 3)
    A univariate time series model can be fitted to each factor and forecasts are generated. The forecast factors can be used to construct forecast functions.
5 Construction of prediction intervals
We equally split the data sample consisting of 48 years from 1975 to 2022 into training, validation, and testing sets, each consisting of 16 years. Using the data in the training sample, we implement an expanding window forecast scheme to obtain the $h$-step-ahead density forecasts in the validation set for $h=1,2,\dots,15$. For each horizon, we have different numbers of curves in the validation set. For instance, when $h=1$, we have 16 years to evaluate the forecast errors, denoted by $\widehat{\epsilon}_{m,s}^{g}(u)=d_{m,s}^{g}(u)-\widehat{d}_{m,s}^{g}(u)$ for $m=1,2,\dots,M=16$; when $h=15$, we have two years to evaluate the residual functions between the samples in the validation set and their corresponding forecasts. From these residual functions, we compute the functional standard deviation, denoted by
$\gamma(u)=\text{sd}[\widehat{\epsilon}_{1,s}^{g}(u),\dots,\widehat{\epsilon}_{M,s}^{g}(u)].$ (8)
We aim to determine $(\overline{\xi}_{\alpha},\underline{\xi}_{\alpha})$ such that $\alpha\times 100\%$ of the residuals satisfy
$-\underline{\xi}_{\alpha}\gamma(u)\leq\widehat{\epsilon}_{m,s}^{g}(u)\leq\overline{\xi}_{\alpha}\gamma(u),$
where $(\overline{\xi}_{\alpha},\underline{\xi}_{\alpha})$ are the tuning parameters.
Typically, the constants $\overline{\xi}_{\alpha}$ and $\underline{\xi}_{\alpha}$ are chosen equal. By a law of large numbers, one may achieve
$\text{Pr}\left[-\xi_{\alpha}\gamma(u)\leq d_{n+h,s}^{g}(u)-\widehat{d}_{n+h,s}^{g}(u)\leq\xi_{\alpha}\gamma(u)\right]\approx\frac{1}{M}\sum^{M}_{m=1}\mathds{1}\left[-\xi_{\alpha}\gamma(u)\leq\widehat{\epsilon}_{m,s}^{g}(u)\leq\xi_{\alpha}\gamma(u)\right],$
where $\mathds{1}(\cdot)$ denotes binary indicator function and $M$ denotes the number of years in the validation set.
To determine the optimal $\xi_{\alpha}$, the samples in the validation set are used to calibrate the prediction intervals so that the empirical coverage probabilities defined in (10) are close to their nominal coverage probabilities.
In Appendix C, we also present the conformal prediction interval approach of Shafer & Vovk (2008) that does not require the tuning parameter. Essentially, by considering a quantile of the absolute residual functions, the pointwise prediction interval for the out-of-sample testing set is centered around the point forecast, with its spread determined by the chosen quantile.
6 Results
6.1 Expanding window scheme
An expanding window analysis of a time-series model is commonly used to assess model and parameter stability and prediction accuracy over time. The expanding window analysis determines the constancy of a model’s parameter by computing parameter estimates and their forecasts over an expanding window through the sample (for details, see Zivot & Wang, 2006, pp. 313–314).
Using the first 32 years from 1975 to 2006 in the Japanese subnational age- and sex-specific life-table death counts, we produce one- to 16-step-ahead forecasts. Through an expanding-window approach, we re-estimate the parameters in the time-series forecasting models using the first 33 years from 1975 to 2007. Forecasts from the estimated models are then produced for one- to 15-step-ahead forecasts. We iterate this process by increasing the sample size by one year until reaching the end of the data period in 2022. This process produces 16 one-step-ahead forecasts, 15 two-step-ahead forecasts, $\dots$, and one 16-step-ahead forecast. We compare these forecasts with the holdout samples to determine the out-of-sample forecast accuracy. In Figure 4, we display a diagram of the expanding window forecast scheme for forecast horizon $h=1$, although we also consider other forecast horizons from $h=2$ to $16$.
TimeTrainFTrainFTrainFTrainF1975:20061975:2007$\vdots$1975:2021Training WindowForecast (F)
Figure 4: A diagram of the expanding-window forecast scheme.
6.2 Point forecast error metrics
Since life-table death counts can be considered a probability density function, we utilize some density evaluation measures. These measures include the discrete version of the KLD (Kullback & Leibler, 1951), the square root of the JSD (Shannon, 1948). For two probability density functions, denoted by $d_{m+\xi}$ and $\widehat{d}_{m+\xi}$, the symmetric version of the KLD, also known as Jeffrey divergence, is defined as
$\text{KLD}(h)=$ $\text{D}_{\text{KL}}(d_{m+\xi}||\widehat{d}_{m+\xi})+\text{D}_{\text{KL}}(\widehat{d}_{m+\xi}||d_{m+\xi})$
=\ = $\frac{1}{111\times(17-h)}\sum^{16}_{\xi=h}\sum_{x=1}^{111}d_{m+\xi}(u_{x})\cdot\left[\ln d_{m+\xi}(u_{x})-\ln\widehat{d}_{m+\xi}(u_{x})\right]+$
$1$×($\frac{1}{111\times(17-h)}\sum^{16}_{\xi=h}\sum_{x=1}^{111}\widehat{d}_{m+\xi}(u_{x})\cdot\left[\ln\widehat{d}_{m+\xi}(u_{x})-\ln d_{m+\xi}(u_{x})\right],$ (9)
which is non-negative, for $h=1,2,\dots,16$.
An alternative metric is the JSD, defined by
$\text{JSD}(h)=\frac{1}{2}D_{\text{KL}}(d_{m+\xi}||\delta_{m+\xi})+\frac{1}{2}D_{\text{KL}}(\widehat{d}_{m+\xi}||\delta_{m+\xi}),$
where $\delta_{m+\xi}$ measures a common quantity between $d_{m+\xi}$ and $\widehat{d}_{m+\xi}$. Following the early work of Shang & Haberman (2020), we consider the geometric mean $\delta_{m+\xi}=\sqrt{d_{m+\xi}\widehat{d}_{m+\xi}}$. To make the JSD a metric between two probability densities, we take the square root of the JSD (see, e.g., Fuglede & Topsoe, 2004).
6.3 Comparison of point forecast accuracy
In Table LABEL:tab:1, averaged across 47 prefectures, we compare the point forecast accuracy among the functional time-series models. From the summary statistics of the KLD and JSD, the HDFPCA model offers the most accurate point forecasts for the female data. For the male data, the MFTS and MLFTS models provide the most accurate point forecasts. For the UFTS, MFTS, MLFTS, and FANOVA models, the number of retained functional principal components was determined by the EVR criterion. For the HDFPCA model, we follow the default setting of Gao et al. (2019), where the number of retained components is six and two in the first and second stages, respectively. In Appendix A, we also provide the results for $K=6$.
Table 1: Averaged across 47 prefectures, we evaluate and compare the point forecast accuracy, measured by the KLD and JSD, among the functional time-series models. The number of components is determined by the EVR criterion.
Female Male
Metric $h$ UFTS MFTS MLFTS FANOVA HDFPCA UFTS MFTS MLFTS FANOVA HDFPCA
KLD 1 0.006 0.010 0.006 0.006 0.008 0.005 0.005 0.004 0.005 0.004
2 0.007 0.011 0.006 0.007 0.010 0.006 0.006 0.005 0.005 0.005
3 0.009 0.013 0.007 0.008 0.010 0.007 0.007 0.005 0.006 0.006
4 0.011 0.014 0.009 0.010 0.011 0.007 0.007 0.006 0.007 0.006
5 0.013 0.016 0.010 0.012 0.012 0.008 0.008 0.007 0.008 0.007
6 0.016 0.017 0.011 0.013 0.010 0.009 0.009 0.008 0.008 0.008
7 0.019 0.019 0.013 0.016 0.015 0.011 0.010 0.008 0.009 0.009
8 0.024 0.021 0.016 0.019 0.012 0.012 0.011 0.010 0.010 0.010
9 0.030 0.024 0.019 0.024 0.018 0.013 0.013 0.011 0.011 0.011
10 0.037 0.028 0.023 0.029 0.011 0.014 0.014 0.012 0.012 0.012
11 0.042 0.033 0.025 0.034 0.024 0.015 0.016 0.011 0.012 0.014
12 0.049 0.027 0.029 0.037 0.009 0.016 0.011 0.011 0.011 0.016
13 0.058 0.032 0.037 0.046 0.031 0.017 0.012 0.014 0.013 0.019
14 0.072 0.039 0.045 0.055 0.011 0.018 0.014 0.016 0.015 0.023
15 0.098 0.047 0.061 0.076 0.045 0.019 0.015 0.021 0.021 0.032
16 0.130 0.059 0.091 0.104 0.015 0.020 0.017 0.031 0.028 0.044
Mean 0.039 0.026 0.026 0.031 0.016 0.012 0.011 0.011 0.011 0.014
Median 0.027 0.023 0.017 0.021 0.011 0.012 0.011 0.010 0.011 0.010
JSD 1 0.038 0.048 0.037 0.044 0.038 0.035 0.034 0.031 0.033 0.034
2 0.042 0.051 0.040 0.046 0.041 0.038 0.036 0.034 0.035 0.036
3 0.045 0.054 0.043 0.049 0.044 0.040 0.039 0.036 0.037 0.038
4 0.051 0.057 0.046 0.048 0.048 0.043 0.041 0.038 0.040 0.040
5 0.056 0.060 0.050 0.056 0.052 0.046 0.043 0.041 0.042 0.042
6 0.061 0.061 0.053 0.050 0.056 0.049 0.045 0.043 0.044 0.043
7 0.066 0.063 0.057 0.062 0.060 0.052 0.047 0.045 0.047 0.045
8 0.073 0.066 0.061 0.053 0.065 0.055 0.049 0.047 0.049 0.047
9 0.081 0.070 0.067 0.070 0.071 0.057 0.052 0.049 0.052 0.049
10 0.089 0.074 0.073 0.054 0.078 0.060 0.054 0.052 0.056 0.051
11 0.099 0.079 0.078 0.080 0.086 0.062 0.056 0.053 0.061 0.052
12 0.109 0.080 0.085 0.053 0.095 0.063 0.054 0.055 0.066 0.053
13 0.121 0.086 0.095 0.093 0.105 0.065 0.057 0.059 0.073 0.057
14 0.134 0.092 0.104 0.056 0.114 0.068 0.060 0.063 0.079 0.060
15 0.159 0.100 0.123 0.112 0.136 0.071 0.062 0.072 0.094 0.070
16 0.188 0.115 0.151 0.066 0.162 0.073 0.065 0.086 0.113 0.083
Mean 0.088 0.072 0.073 0.062 0.078 0.055 0.050 0.050 0.058 0.050
Median 0.077 0.068 0.064 0.055 0.068 0.056 0.050 0.048 0.051 0.050
Figure 5 presents two heatmaps comparing seven functional time-series models across 16 forecast horizons for both female and male data. For each horizon, we count how often each method produces the smallest point forecast errors, as measured by the KLD. Choosing $K=6$ has a slight advantage compared to the one based on the EVR criterion. For shorter forecast horizons, the UFTS, MFTS, and MLFTS models perform equally well for the female data. However, for relatively longer forecast horizons, the HDFPCA model is recommended. In the case of male data, the MLFTS model is generally preferred; only at the shorter forecast horizon is $K=6$ advantageous. For relatively longer forecast horizons, the FANOVA with the EVR criterion is recommended.
(a) Female data
(b) Male data
Figure 5: Heatmaps comparing the point forecast accuracy, measured by the KLD, between nine functional time-series models across 16 forecast horizons for the testing female and male data. Since there are 47 prefectures, each row sums to 47. A method appearing more than five times is highlighted in green, while those appearing fewer than five times are shown in red. If a method appears exactly five times, it is displayed in white.
6.4 Interval forecast error metrics
To evaluate interval forecast accuracy, we consider the CPD between the empirical coverage probability (ECP) and nominal coverage probability, as well as the mean interval score of Gneiting & Raftery (2007). For each year in the forecasting period, the $h$-step-ahead prediction intervals are computed at the $100(1-\alpha)\%$ nominal coverage probability. We consider the common case of the symmetric $100(1-\alpha)\%$ prediction intervals, with lower and upper bounds that are quantiles at $\alpha/2$ and $21-\alpha/$ 2, denoted by $\widehat{d}_{n+\xi,s}^{g,\text{lb}}(u)$ and $\widehat{d}_{n+\xi,s}^{g,\text{ub}}(u)$. The ECP and CPD are defined as
$\text{ECP}_{h}$ $=\frac{1}{111\times(16-h)}\times\sum_{\xi=h}^{15}\sum_{u=1}^{111}\mathds{1}\left\{\widehat{d}_{\nu+\xi,s}^{g,\text{lb}}(u)\leq d_{\nu+\xi,s}^{g}(u)\leq\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u)\right\},$ (10)
$\text{CPD}_{h}$ $=\frac{1}{111\times(16-h)}\times\sum_{\xi=h}^{15}\sum_{u=1}^{111}\left[\mathds{1}\{d_{\nu+\xi,s}^{g}(u)>\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u)\}+\mathds{1}\{d_{\nu+\xi,s}^{g}(u)<\widehat{d}_{\nu+\xi,s}^{g,\text{lb}}(u)\}\right],$
where $h=1,\dots,15$, $v$ denotes the end of training set when we calibrate the CPD or interval score, and $\nu$ denotes the end of validation set when we evaluate the interval forecast accuracy.
For different ages and years in the testing set, the mean and median CPD are defined as
$\overline{\text{CPD}}$ $=\frac{1}{15}\sum^{15}_{h=1}\text{CPD}_{h},$
M[CPD] $=\text{median}(\text{CPD}_{h}).$
As defined by Gneiting & Raftery (2007), a scoring rule for the prediction intervals at age $u$ is
$S_{\alpha,\xi}\left[\widehat{d}^{g,\text{lb}}_{\nu+\xi,s}(u),\widehat{d}^{g,\text{ub}}_{\nu+\xi,s}(u),d^{g}_{\nu+\xi,s}(u)\right]$ $=\left[\widehat{d}^{g,\text{ub}}_{\nu+\xi,s}(u)-\widehat{d}^{g,\text{lb}}_{\nu+\xi,s}(u)\right]$
$+\frac{2}{\alpha}\left[\widehat{d}^{g,\text{lb}}_{\nu+\xi,s}(u)-d_{n+\xi,s}^{g}(u)\right]\mathds{1}\left\{d_{\nu+\xi,s}^{g}(u)<\widehat{d}_{\nu+\xi,s}^{g,\text{lb}}(u)\right\}$
$+\frac{2}{\alpha}\left[d_{\nu+\xi,s}^{g}(u)-\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u)\right]\mathds{1}\left\{d_{\nu+\xi,s}^{g}(u)>\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u)\right\},$
where the level of significance is customarily set to be $\alpha=0.2$ or 0.05. The interval score rewards a narrow prediction interval width if and only if $100(1-\alpha)\%$ of the holdout densities lie within the prediction interval.
For different ages and years in the testing set, the mean interval score is defined by
$\overline{S}_{\alpha}(h)=\frac{1}{111\times(16-h)}\times\sum^{15}_{\xi=h}\sum^{111}_{u=1}S_{\alpha,\xi}\left[\widehat{d}_{\nu+\xi,s}^{g,\text{lb}}(u),\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u),d_{\nu+\xi,s}^{g}(u)\right].$
Averaging over all forecast horizons, we obtain the overall mean interval score
$\overline{S}_{\alpha}$ $=\frac{1}{15}\sum^{15}_{h=1}\overline{S}_{\alpha}(h),$
$\text{M}[S_{\alpha}]$ $=\text{median}[\overline{S}_{\alpha}(h)].$
6.5 Comparison of interval forecast accuracy
In Table LABEL:tab:2, averaged across 47 prefectures, we compare the interval forecast accuracy among the functional time-series models using the EVR criterion. From the summary statistics of the CPD and interval score at the level of significance $\alpha=0.2$, the MFTS offers the smallest CPD value for the female data, while the UFTS and MLFTS offer the smallest CPD values for the male data. In terms of the interval scores, the HDFPCA provides the smallest interval scores for the female data, while the MLFTS gives the smallest interval scores for the male data. In Appendix B, we also provide the results for $K=6$.
Table 2: Averaged across 47 prefectures, we evaluate and compare the interval forecast accuracy, as measured by the CPD and interval score, among the seven functional time-series models at the level of significance $\alpha=0.2$. The number of components is determined by the EVR criterion.
Female Male
Metric $h$ UFTS MFTS MLFTS FANOVA HDFPCA UFTS MFTS MLFTS FANOVA HDFPCA
$\overline{\text{CPD}}_{0.2}$ 1 0.061 0.065 0.043 0.075 0.072 0.056 0.085 0.044 0.074 0.059
2 0.067 0.058 0.045 0.088 0.053 0.058 0.088 0.047 0.079 0.071
3 0.071 0.054 0.052 0.097 0.074 0.055 0.090 0.052 0.088 0.073
4 0.081 0.051 0.058 0.117 0.076 0.054 0.094 0.057 0.093 0.084
5 0.087 0.048 0.065 0.122 0.091 0.056 0.096 0.061 0.098 0.086
6 0.086 0.045 0.069 0.128 0.100 0.058 0.095 0.058 0.095 0.090
7 0.090 0.047 0.071 0.130 0.069 0.057 0.091 0.056 0.087 0.084
8 0.095 0.046 0.074 0.131 0.107 0.060 0.085 0.056 0.089 0.075
9 0.096 0.050 0.080 0.131 0.060 0.072 0.083 0.060 0.087 0.076
10 0.104 0.052 0.088 0.143 0.138 0.073 0.082 0.063 0.087 0.086
11 0.103 0.051 0.082 0.144 0.048 0.075 0.086 0.074 0.084 0.084
12 0.105 0.059 0.087 0.136 0.134 0.069 0.083 0.067 0.080 0.077
13 0.106 0.061 0.093 0.137 0.048 0.069 0.082 0.080 0.080 0.080
14 0.116 0.067 0.094 0.145 0.096 0.065 0.079 0.076 0.089 0.077
15 0.122 0.069 0.093 0.137 0.086 0.057 0.067 0.071 0.070 0.085
Mean 0.093 0.055 0.073 0.118 0.083 0.062 0.086 0.062 0.085 0.079
Median 0.095 0.052 0.074 0.131 0.076 0.058 0.085 0.060 0.087 0.080
$\overline{S}_{0.2}$ 1 255 393 246 279 314 271 270 213 263 243
2 299 417 279 328 344 298 296 237 283 270
3 345 445 312 384 369 325 322 259 300 292
4 412 476 353 470 361 356 351 286 321 321
5 477 511 397 551 464 390 376 309 345 349
6 549 543 429 639 381 419 397 331 358 370
7 626 565 476 734 537 445 412 353 375 387
8 721 600 536 843 420 465 433 378 393 409
9 829 650 606 973 625 493 451 402 416 436
10 956 725 690 1133 476 525 481 432 439 488
11 1070 798 767 1271 793 543 506 430 455 519
12 1199 788 864 1429 500 570 467 451 460 560
13 1350 874 1000 1623 966 583 492 507 510 623
14 1501 945 1108 1823 546 661 547 582 571 669
15 1888 1122 1419 2224 1258 875 690 811 835 933
Mean 832 657 632 1000 557 481 433 399 420 458
Median 721 600 536 843 476 465 433 378 393 409
In Figure 6, we present two heatmaps comparing seven functional time-series models across 15 forecast horizons for both female and male testing data. For each horizon, we count how often each method produces the smallest interval forecast errors, as measured by the CPD and interval score. By examining the CPD results, we recommend the MFTS for the female data and the MLFTS model for the male data, especially for $K=6$. By examining the interval score results, we recommend the MFTS for shorter horizons and the HDFPCA for longer horizons in the female data. For the male data, the MLFTS generally provides the smallest interval scores.
(a) CPD for females
(b) CPD for males
(c) Interval score for females
(d) Interval score for males
Figure 6: Heatmaps comparing the interval forecast accuracy at the level of significance $\alpha=0.2$, measured by the CPD and interval score, between seven functional time-series models across 15 forecast horizons for the female and male testing data. Since there are 47 prefectures, each row sums to 47. A method appearing more than five times is shown in green, while those appearing fewer than five times are shown in red. If a method appears exactly five times, it is shown in white.
The interval forecast results at the significance level of $\alpha=0.05$ are available and can be viewed in a developed Shiny app at https://cfjimenezv.shinyapps.io/Shiny_App_CDF/.
7 Conclusion
We present several graphical tools, namely image plots and cross-correlation plots, to visualize the differences and dependence in life-table death counts between each prefecture and Japan as a whole.
Via the CDF transformation, we consider a suite of functional time-series models to model and forecast subnational age distribution of death counts for 47 Japanese prefectures. The CDF transformation is ideal to handle the zero values in the subnational data. Within the CDF transformation, we also present two general strategies for constructing pointwise prediction intervals.
By dividing the datasets into training and testing samples, we evaluate and compare point and interval forecast accuracy among the functional time-series models. We confirm the importance of joint modeling techniques, in which the MFTS, MLFTS, FANOVA, and HDFPCA models generally perform well on the aggregate level across 47 prefectures. We present a novel heatmap to show the frequency (of 47 prefectures), in which one method performs the best for each forecast horizon. The individual forecast errors for various horizons, obtained from all methods for each prefecture, are available in a developed shiny app at https://cfjimenezv.shinyapps.io/Shiny_App_CDF/.
There are several ways the methodology can be further extended, and we briefly list four. First, the functional standard deviation in (8) was computed coordinate-wise. A range of functional depth exists, which can be implemented to compute other variants of standard deviations. Second, instead of symmetric prediction intervals, asymmetric ones can be considered, searching for two tuning parameters to adjust the lower and upper bounds. Third, the data set was divided equally into the training, validation, and testing samples. Other proportions may be possible and lead to a more accurate selection of tuning parameter $\xi_{\alpha}$ and more accurate interval forecasts. Fourth, to demonstrate we implemented a suite of functional time-series models. Other time-series extrapolation models may also be considered, such as the robust functional principal component analysis by Oguamalam et al. (2025). It extends Mahalanobis distance to Bayes spaces and introduces a regularized approach for improved principal component estimation in the presence of outliers.
Supplementary Materials
  1. 1.
    Code for the forecasting methods based on the CDF transformation.The [Uncaptioned image] code to produce point and interval forecasts from the different functional time-series models described in the paper. The [Uncaptioned image] codes are available at the following repository: https://github.com/cfjimenezv07/HDFTS_CDF/tree/main/Rcode
  2. 2.
    Code for shiny application.The [Uncaptioned image] code to produce a shiny user interface for the point and interval forecasts for the 47 Japanese prefectures. The [Uncaptioned image] codes are available at the following repository: https://github.com/cfjimenezv07/HDFTS_CDF/tree/main/Shiny_App_CDF
Acknowledgment
This research was supported by the Australian Research Council Discovery Project (grant no. DP230102250) and Future Fellowship (grant no. FT240100338). The author is grateful to Dr Yang Yang from the University of Newcastle for assistance with the heatmaps.
Appendix A: Point forecast accuracy when $K=6$
In Table LABEL:tab:appendix_1, averaged across 47 prefectures, we compare the point forecast accuracy among the functional time-series models, when the number of retained components is set to be six as in Hyndman et al. (2013). From the summary statistics of the KLD and JSD, the MFTS offers the most accurate point forecasts for the female and male data.
Table 3: Averaged across 47 prefectures, we evaluate and compare the point forecast accuracy, measured by the KLD and JSD, among the functional time-series models. The number of components is determined by setting $K=6$.
Female Male
Metric $h$ UFTS MFTS MLFTS FANOVA UFTS MFTS MLFTS FANOVA
KLD 1 0.005 0.006 0.005 0.005 0.004 0.004 0.004 0.004
2 0.007 0.007 0.006 0.006 0.005 0.005 0.004 0.005
3 0.008 0.008 0.007 0.008 0.006 0.005 0.005 0.005
4 0.010 0.010 0.009 0.009 0.007 0.006 0.006 0.006
5 0.013 0.012 0.011 0.011 0.008 0.007 0.007 0.007
6 0.015 0.013 0.012 0.013 0.009 0.008 0.008 0.008
7 0.018 0.015 0.014 0.016 0.010 0.009 0.009 0.009
8 0.023 0.017 0.017 0.019 0.011 0.010 0.010 0.010
9 0.029 0.021 0.021 0.024 0.012 0.011 0.011 0.012
10 0.035 0.025 0.025 0.029 0.013 0.013 0.012 0.013
11 0.041 0.029 0.028 0.034 0.014 0.014 0.011 0.013
12 0.048 0.026 0.033 0.039 0.015 0.010 0.012 0.012
13 0.057 0.032 0.042 0.048 0.016 0.011 0.014 0.014
14 0.070 0.040 0.049 0.057 0.018 0.013 0.016 0.015
15 0.096 0.050 0.068 0.079 0.019 0.014 0.020 0.021
16 0.128 0.067 0.097 0.108 0.020 0.016 0.029 0.028
Mean 0.038
0.024
0.028 0.032 0.012
0.010
0.011 0.011
Median 0.026
0.019
0.019
0.022 0.012
0.010
0.011 0.011
JSD 1 0.037 0.038 0.036 0.036 0.033 0.031 0.030 0.031
2 0.040 0.041 0.039 0.039 0.036 0.033 0.033 0.034
3 0.044 0.044 0.042 0.042 0.038 0.035 0.035 0.036
4 0.049 0.047 0.046 0.047 0.041 0.038 0.038 0.038
5 0.054 0.051 0.050 0.051 0.044 0.040 0.040 0.041
6 0.059 0.054 0.054 0.055 0.047 0.042 0.043 0.043
7 0.065 0.057 0.058 0.060 0.050 0.044 0.045 0.045
8 0.072 0.060 0.063 0.065 0.054 0.047 0.048 0.048
9 0.079 0.065 0.069 0.072 0.056 0.049 0.050 0.050
10 0.088 0.071 0.076 0.080 0.059 0.052 0.053 0.053
11 0.097 0.076 0.083 0.088 0.061 0.054 0.054 0.054
12 0.108 0.078 0.091 0.097 0.063 0.052 0.056 0.056
13 0.120 0.086 0.101 0.107 0.065 0.055 0.060 0.059
14 0.133 0.093 0.110 0.117 0.068 0.059 0.064 0.063
15 0.157 0.104 0.131 0.140 0.070 0.061 0.071 0.072
16 0.186 0.123 0.158 0.168 0.073 0.064 0.085 0.084
Mean 0.087
0.068
0.076 0.079 0.054
0.047
0.050 0.050
Median 0.075
0.063
0.066 0.068 0.055
0.048
0.049 0.049
Appendix B: Interval forecast accuracy when $K=6$
In Table LABEL:tab:appendix_2, averaged across 47 prefectures, we compare the interval forecast accuracy among the functional time-series models. From the summary statistics of the CPD and interval scores at the level of significance $\alpha=0.2$, the MFTS offers the smallest CPD value for the female data, while the MLFTS offers the smallest CPD values for the male data. In terms of the interval scores, the MFTS provides the smallest interval scores for the female data, while the MFTS and MLFTS present about the same and equally smallest interval scores for the male data.
Table 4: Averaged across 47 prefectures, we evaluate and compare the interval forecast accuracy, as measured by the CPD and interval score, among the functional time-series models at the level of significance $\alpha=0.2$. The number of components is determined by setting $K=6$. For either females or males, we highlight the method with the smallest overall CPD and interval scores.
Female Male
Metric $h$ UFTS MFTS MLFTS FANOVA UFTS MFTS MLFTS FANOVA
CPD 1 0.054 0.040 0.046 0.056 0.052 0.043 0.043 0.049
2 0.064 0.042 0.051 0.060 0.054 0.051 0.043 0.050
3 0.068 0.045 0.059 0.068 0.056 0.056 0.047 0.054
4 0.076 0.049 0.064 0.083 0.056 0.065 0.049 0.058
5 0.083 0.051 0.069 0.092 0.061 0.073 0.056 0.066
6 0.086 0.052 0.071 0.099 0.059 0.076 0.052 0.073
7 0.087 0.053 0.072 0.105 0.059 0.079 0.052 0.067
8 0.091 0.056 0.071 0.112 0.063 0.081 0.055 0.074
9 0.096 0.056 0.075 0.111 0.073 0.080 0.059 0.078
10 0.104 0.056 0.082 0.125 0.073 0.080 0.061 0.081
11 0.103 0.050 0.078 0.122 0.074 0.085 0.068 0.085
12 0.106 0.058 0.082 0.113 0.065 0.084 0.067 0.076
13 0.106 0.065 0.084 0.113 0.067 0.078 0.069 0.078
14 0.116 0.068 0.091 0.122 0.066 0.077 0.069 0.074
15 0.121 0.068 0.086 0.122 0.056 0.069 0.064 0.065
Mean 0.091
0.054
0.072 0.100 0.062 0.072
0.057
0.068
Median 0.091
0.053
0.072 0.111 0.061 0.077
0.056
0.073
$\overline{S}_{0.2}$ 1 232 244 227 242 242 219 204 229
2 278 275 265 294 271 247 230 254
3 324 309 301 348 302 275 257 281
4 388 347 349 428 333 303 287 309
5 454 392 396 504 370 334 315 341
6 525 438 435 587 402 361 341 368
7 599 477 486 678 431 378 367 393
8 693 526 552 785 454 401 393 418
9 798 587 625 915 482 422 421 444
10 925 673 716 1068 513 453 453 478
11 1041 747 799 1210 529 482 448 495
12 1174 764 902 1364 559 452 465 512
13 1329 869 1042 1557 574 484 514 560
14 1474 956 1159 1753 644 540 590 633
15 1863 1145 1445 2124 865 698 819 868
Mean 806
583
647 924 465
403
407 439
Median 693
526
552 785 454 401
393
418
Appendix C: Conformal prediction intervals
In machine learning, a popular methodology known as conformal prediction (Shafer & Vovk, 2008) is used to construct probabilistic forecasts calibrated on out-of-sample errors. From the absolute value of $[\widehat{\varepsilon}_{1,s}^{g}(u),\dots,\widehat{\varepsilon}_{M,s}^{g}(u)]$, we compute its $100(1-\alpha)\%$ quantile for a level of significance $\alpha$, denoted by $q_{\alpha}(u)$. The prediction interval can be obtained as
$\left[\widehat{d}^{g}_{n+h|n,s}(u)-q_{\alpha}(u),\quad\widehat{d}^{g}_{n+h|n,s}(u)+q_{\alpha}(u)\right],$
where $\widehat{d}^{g}_{n+h|n,s}(u)$ denotes the $h$-step-ahead point forecasts for the data in the testing set.
The conformal prediction works well when the sample size in the validation set is reasonably large (Dhillon et al., 2024). For a small set of samples, we suggest the standard deviation approach in Section 5. In Table LABEL:tab:appendix_3, the MFTS provides the smallest CPD values under the EVR criterion, while the MLFTS offers the smallest CPD values under $K=6$ for the female data. For the male data, the MLFTS is recommended. In terms of interval scores, the MFTS and MLFTS perform on par and are equally the best for the female data. For the male data, the MLFTS is preferable.
Table 5: Averaged across 47 prefectures, we evaluate and compare the interval forecast accuracy, measured by the CPD and interval score, among the functional time-series models at the level of significance $\alpha=0.2$. The number of components is determined by the EVR criterion or $K=6$; for each we highlight the method with the smallest overall CPD and interval score values.
EVR $K=6$
Metric Sex $h$ UFTS MFTS MLFTS FANOVA HDFPCA UFTS MFTS MLFTS FANOVA
CPD F 1 0.080 0.096 0.055 0.096 0.098 0.067 0.051 0.047 0.067
2 0.087 0.087 0.054 0.110 0.049 0.075 0.045 0.052 0.079
3 0.093 0.082 0.058 0.124 0.105 0.083 0.052 0.057 0.091
4 0.105 0.078 0.062 0.151 0.043 0.096 0.054 0.064 0.110
5 0.116 0.077 0.066 0.161 0.128 0.108 0.059 0.067 0.123
6 0.124 0.073 0.063 0.172 0.047 0.118 0.061 0.065 0.136
7 0.129 0.066 0.066 0.170 0.122 0.123 0.058 0.068 0.140
8 0.133 0.063 0.066 0.172 0.056 0.130 0.058 0.068 0.145
9 0.138 0.063 0.071 0.177 0.121 0.136 0.063 0.072 0.156
10 0.147 0.074 0.082 0.186 0.049 0.147 0.075 0.083 0.171
11 0.151 0.083 0.082 0.192 0.138 0.150 0.086 0.091 0.172
12 0.148 0.085 0.090 0.186 0.052 0.150 0.090 0.099 0.166
13 0.151 0.093 0.099 0.193 0.151 0.150 0.098 0.107 0.167
14 0.171 0.121 0.124 0.212 0.212 0.169 0.122 0.132 0.185
15 0.214 0.159 0.165 0.263 0.263 0.208 0.165 0.173 0.222
Mean 0.132 0.087
0.080
0.171 0.098 0.127
0.076
0.083 0.142
Median 0.133 0.082
0.066
0.172 0.098 0.130
0.061
0.068 0.145
M 1 0.076 0.109 0.056 0.094 0.083 0.066 0.059 0.055 0.058
2 0.078 0.112 0.063 0.104 0.096 0.073 0.069 0.058 0.066
3 0.078 0.114 0.069 0.109 0.101 0.074 0.076 0.065 0.072
4 0.082 0.124 0.086 0.121 0.119 0.079 0.087 0.074 0.082
5 0.088 0.131 0.095 0.129 0.124 0.090 0.098 0.086 0.099
6 0.098 0.140 0.105 0.134 0.137 0.102 0.114 0.096 0.115
7 0.096 0.131 0.096 0.125 0.134 0.106 0.113 0.091 0.119
8 0.099 0.130 0.098 0.125 0.127 0.110 0.113 0.093 0.126
9 0.107 0.132 0.095 0.127 0.134 0.120 0.120 0.091 0.134
10 0.114 0.140 0.098 0.128 0.153 0.125 0.133 0.099 0.137
11 0.106 0.147 0.095 0.122 0.164 0.121 0.139 0.096 0.141
12 0.108 0.142 0.092 0.115 0.162 0.118 0.140 0.099 0.139
13 0.107 0.142 0.090 0.106 0.162 0.119 0.142 0.096 0.138
14 0.117 0.163 0.097 0.124 0.172 0.135 0.160 0.106 0.156
15 0.144 0.192 0.142 0.149 0.196 0.161 0.192 0.142 0.181
Mean 0.100 0.137
0.092
0.121 0.138 0.107 0.117
0.090
0.127
Median 0.099 0.132
0.095
0.124 0.134 0.110 0.114
0.093
0.116
$\overline{S}_{0.2}$ F 1 255 386 241 279 323 235 249 230 248
2 302 412 277 329 329 283 282 269 301
3 351 442 312 387 379 331 318 306 358
4 421 473 358 476 352 400 356 356 441
5 489 504 407 553 467 468 400 408 517
6 565 526 446 642 377 543 439 451 603
7 642 554 494 734 527 618 477 502 695
8 741 581 551 846 432 714 516 563 807
9 853 628 620 979 609 826 570 639 941
10 986 686 699 1147 456 961 638 725 1104
11 1106 756 766 1287 745 1084 705 801 1249
12 1242 734 859 1443 467 1220 713 904 1415
13 1383 822 994 1639 899 1361 817 1035 1608
14 1565 909 1098 1832 490 1541 920 1152 1802
15 1887 1018 1321 2184 1181 1852 1054 1395 2142
Mean 853 629 630 984
535
829 564 649 949
Median 741 581 551 846
467
714 516 563 807
M 1 258 273 218 266 247 233 220 205 229
2 282 297 241 287 275 260 247 230 253
3 306 323 262 301 299 287 274 256 278
4 329 350 289 321 331 313 301 285 302
5 356 373 311 342 360 344 330 311 331
6 377 391 329 354 385 369 353 334 355
7 397 407 346 368 408 394 372 356 377
8 418 425 369 384 429 415 391 380 404
9 444 443 389 403 463 444 411 404 429
10 462 470 414 422 513 462 443 428 455
11 484 495 408 440 561 484 468 424 479
12 490 453 422 443 596 493 436 436 490
13 509 482 471 474 660 509 467 486 525
14 560 524 521 510 716 555 509 533 567
15 618 560 625 586 874 609 555 624 631
Mean 419 418
374
393 475 411 385
379
407
Median 418 425
369
348 429 415 391
380
404