### v2

# Modeling and forecasting subnational age distribution of death counts

Abstract
Existing mortality forecasting methods focus on age-specific mortality rates, which lie in an unconstrained space and overlook the distributional nature of life-table death counts. Few studies have developed and compared forecasting methods that model the shape and dynamics of the age distribution of deaths, especially at the subnational level, where data quality varies greatly. This paper presents several forecasting methods to model and forecast the subnational age distribution of death counts. The age distribution of death counts has many similarities to probability density functions, which are nonnegative and have a constrained integral, and thus live in a constrained nonlinear space. To address the nonlinear nature of objects, we implement a cumulative distribution function transformation that is scale-free and has additional monotonicity. Using subnational Japanese life-table death counts from Japanese Mortality Database (2025), we evaluate the forecast accuracy of the transformation and forecasting methods. The improved forecast accuracy of life-table death counts implemented here will be of great interest to demographers in estimating regional age-specific survival probabilities and life expectancy, and to actuaries for determining annuity prices for various ages and maturities.
Keywords: conformal prediction intervals; constrained functional time series; cumulative distribution function transformation; life-table death counts; principal component analysis
1 Introduction
Accurate prediction of subnational mortality rates plays a crucial role in both actuarial studies and demographic planning. From a demographic perspective, subnational mortality forecasts inform regional health planning, pension sustainability assessments, and population projections that guide policy decisions on healthcare infrastructure and resource allocation (see, e.g., Booth & Tickle, 2008). In actuarial science, understanding spatial variations in mortality and their evolution over time is essential for pricing life insurance products, designing annuities, and managing longevity risk in pension schemes (e.g., Cairns et al., 2006; Haberman & Renshaw, 2011). Subnational mortality prediction allows actuaries and policymakers to capture heterogeneity across regions, reflecting socioeconomic, environmental, and healthcare disparities, improving the reliability of longevity projections (Li & Lee, 2005).
Although there exist many multipopulation techniques, such as the Li & Lee’s (2005) method, for modeling and forecasting age-specific mortality rates, the age distribution of death counts is another mortality instrument, which possesses its unique advantages over mortality rates. For example, the age distribution of death counts preserves distributional information: the age distribution of deaths retains the shape of how deaths are spread across ages. Due to non-negativity and summability, the changes are relative in age. It captures shifts in longevity more directly: the mode and spread of the distribution can be measured. It is less sensitive to population size estimates; at the subnational level, these estimates may be unreliable. Although mortality rates at older ages can be outside their plausible range (0,1), it does not occur for the age distribution of death counts. It facilitates comparison across populations and time, as it is insensitive to population size estimates. The age distribution of death counts is an explicit way of measuring the probability of dying, commonly studied in the Cairns et al.’s (2006) model.
Following the early works of Lu et al. (2020 and 2021) on subnational mortality rates and Bergeron-Boucher et al. (2018) on subnational age distribution of death counts, we present several forecasting methods to model and predict the subnational age distribution of deaths, which are also useful for applications in actuarial science, such as fixed-term or lifetime annuity pricing (see, e.g., Shang & Haberman, 2020). Apart from subnational age distributions of deaths in demography (Jiménez-Varón et al., 2025), density-valued objects for multiple groups are common, with examples including income distributions across different populations (Kneip & Utikal, 2001), financial return distributions for multiple stocks (Petersen et al., 2022), distributions of bidding times in online auctions for various items (Wang et al., 2008), among others.
Since age distributions of deaths naturally take the form of density functions, they are well-suited for modeling using advanced forecasting techniques from compositional and functional data analyses. Oeppen (2008) demonstrates that the use of compositional data analysis to forecast the age distribution of death counts performs similarly in accuracy to forecast age-specific mortality rates. Within compositional data analysis, the Lee-Carter model is frequently applied to model and forecast unconstrained data. Shang & Haberman (2020) extend this approach by incorporating a functional data-analytic perspective, including multiple principal components and nonparametric smoothing. Recently, Shang & Haberman (2025b) introduce a cumulative distribution function (CDF) transformation, converting each year’s age distribution of deaths into a probability bounded within the unit interval to address the presence of zero values. Through cumulative summation, a probability density function can be transformed into a CDF, providing the additional benefit of monotonicity (see also Mayhew & Smith, 2013). With a time series of CDFs, it is common to model and extrapolate the pattern using a logistic transformation.
Multiple density-valued objects observed over time are related to high-dimensional functional time series (HDFTS). In the statistical literature, Zhou & Dette (2023) derive Gaussian and multiplier bootstrap approximations for the sums of HDFTS. With these approximations, they construct joint simultaneous confidence bands for the mean functions and develop a hypothesis to test whether the mean functions in the cross-sectional dimension exhibit parallel behavior. Hallin et al. (2023) investigate the representation of HDFTS using a factor model and determine conditions on the eigenvalues of the covariance operator that are crucial for establishing the existence and uniqueness of the factor model. Gao et al. (2019) adopt a two-stage approach, combining truncated principal component analysis and a separate scalar factor model for the resulting panels of scores. Tavakoli et al. (2023) introduce a functional factor model with a functional factor loading and a vector of real-valued factors, while Guo et al. (2025) propose a functional factor model with a real-valued factor loading and a functional factor. Leng et al. (2026) introduce a unified framework that accommodates both types of factor models. Tang et al. (2022) study clustering for age-specific subnational mortality rates, which is an example of HDFTS. Li et al. (2024) introduce hypothesis tests and estimation procedures for testing and locating (common) change points. Chang et al. (2025) develop a two-stage procedure for modeling and forecasting HDFTS.
Although numerous studies have addressed the modelling and forecasting of age distributions of death counts at the national level (see, e.g., Shang, 2025; Shang & Haberman, 2025a and b), there remains a notable gap in extending such forecasts to the subnational level. Our contributions are twofold. First, we present several visualization techniques to capture patterns in the age distribution of death counts and to display contrasts between subnational and national data. Second, we revisit several forecasting methods and compare their point and interval forecast accuracy with applications to subnational life-table death counts. For different sexes and prefectures, we provide some recommendations on the optimal multipopulation models, which can vary in terms of their point and interval forecast accuracies. The improvement in forecast accuracy of life-table death counts may translate to more accurate annuity prices and life expectancy calculations; in turn, these inform policymakers on the sustainability of the pension system and aged care facilities.
Our paper is structured as follows. In Section 2, we describe our motivating data set and introduce a series of image plots for displaying important features in HDFTS. In Section 3, we present the CDF transformations. Within the CDF transformation, we consider a number of forecasting methods in Section 4 to model and forecast unconstrained functional time-series data. In Section 5, we propose two general strategies for constructing pointwise prediction intervals for the age distribution of death counts. In Section 6, we evaluate and compare point forecast accuracy using the Kullback-Leibler divergence (KLD) and Jensen-Shannon divergence (JSD), and interval forecast accuracy using the coverage probability difference (CPD) and mean interval score. In Section 6.6, we show relationship between life-table death count and survival probability and life expectancy. In Section 6.7, we present a model diagnostic analysis using functional analysis of variance. Section 7 concludes with ideas on how the methodology presented can be extended.
2 Subnational age distribution of death in Japan
In many developed countries, such as Japan, increases in longevity and an aging population have led to concerns about the sustainability of pension, health, and aged care systems (see, e.g., Coulmas, 2007). These concerns have led to a surge of interest among government policymakers in accurately modeling and forecasting age-specific mortality. Subnational forecasts of age-specific mortality are useful for informing policy at the local level, and any improvement in forecast accuracy is beneficial for allocating current and future resources at national and subnational levels.
In demography, the age distribution of deaths provides important insights into longevity and lifespan variability that cannot be grasped directly from the central mortality rate or survival function. As pointed out by Oeppen (2008), the age distribution of death counts is more suitable than central mortality rates for computing life expectancy and annuity premia. In addition to providing an informative description of the mortality experience of a population, life-table death counts yield readily available information on the “central longevity indicators” (e.g., mean, median, and modal age at death, see Cheung et al., 2005; Canudas-Romo, 2010) and lifespan variability (e.g., van Raalte & Caswell, 2013; Aburto & van Raalte, 2018). The life-table death count is less sensitive to population size estimate, especially at the subnational level; population size estimate may be unreliable. Because of this, it is insensitive to population size estimates, which facilitates comparisons across populations and over time. The life-table death count is directly linked to the probability of dying $q_{x}$, which is the mortality instrument in the Cairns et al.’s (2006) model.
We study Japanese life-table death counts from 1975 to 2023, obtained from the Japanese Mortality Database (2025). We observe complete life-table death counts for ages from 0 to 109 in single years of age, with the last age group being age 110+. A distinctive advantage of life-table death counts lies in the availability of data observed across all ages.
(a) Japanese female data
(b) Japanese male data
Figure 1: Functional time series graphical displays of age-specific life-table death counts from 1975 to 2023 in a single-year group. The oldest years are shown in red, with the most recent years in violet. Curves are ordered chronologically according to the colors of the rainbow.
Figure 1 shows rainbow plots of the female and male age-specific life-table death counts in Japan from 1975 to 2023. The time ordering of the curves follows the color order of a rainbow, where curves from the distant past are shown in red, and the more recent curves are shown in violet. The figures show typical mortality curves for a developed country, with a declining trend in infant mortality. A typical negatively skewed distribution of life-table death counts is apparent, with the peaks shifting to higher ages for both females and males. This shift is a major source of longevity risk, which is a major issue for insurers and pension funds, especially in the sale and risk management of annuity products (see Denuit et al., 2007, for a discussion). In addition, the spread of the distribution indicates the variability in lifespan. A decrease in variability over time can be directly observed and quantified using the Gini coefficient (Wilmoth & Horiuchi, 1999; Debón et al., 2017).
In Figure 2, we present a map of 47 Japanese prefectures from the northern region, Hokkaido, to the southern region, Okinawa. Due to geography and socio-economic status, the age distributions of death counts tend to be heterogeneous. Differences between prefectures or years within the same prefecture can be visualized using image plots in Section 2.1 and quantified using the functional cross-correlation function in Section 2.2.
Figure 2: A geographical map of 47 prefectures in Japan.
2.1 Image plots
Another visual perspective of the data is shown in the image plot of Figure 3. To measure the distance between two probability densities, such as the subnational and national age distributions of death counts, we consider the symmetric KLD (see Edwards & Tuljapurkar, 2005; d’Albis et al., 2014, for examples of using the KLD to compare country-specific lifespan distributions). For a given prefecture and gender, the symmetric KLD can be expressed as
$KLD_{s}^{g}=\$$D_{KL}\left[d_{t,s}^{g}(u)|d_{t,\text{national}}^{g}(u)\right]+D_{\text{KL}}\left[d_{t,\text{national}}^{g}(u)|d_{t,s}^{g}(u)\right]$
$=\$$\frac{1}{111\times n}\sum^{n}_{t=1}\sum_{i=1}^{111}d_{t,s}^{g}(u_{i})\times[\ln d_{t,s}^{g}(u_{i})-\ln d_{t,\text{national}}^{g}(u_{i})]+$
$\frac{1}{111\times n}\sum^{n}_{t=1}\sum_{i=1}^{111}d_{t,\text{national}}^{g}(u_{i})\times[\ln d_{t,\text{national}}^{g}(u_{i})-\ln d_{t,s}^{g}(u_{i})],$
where $n$ denotes the number of years in the data set. Here, we graph the symmetric KLD between life-table death counts for each prefecture to life-table death counts for the whole country, thus allowing relative mortality comparisons to be made.
A divergent color palette is used, with red representing larger KLD values (i.e., greater difference between subnational and national data) and yellow denoting smaller KLD values. The prefectures are ordered geographically from north to south, so the most northerly prefecture (Hokkaido) is at the top of the panels, and the most southerly prefecture (Okinawa) is at the bottom.
Figure 3: Image plots showing the symmetric KLD between life-table death counts of a prefecture and life-table death counts of the whole country. The top panel shows mortality averaged over ages, while the bottom panel shows mortality averaged over years. Prefectures are numbered geographically from north to south.
The top row of the panels shows mortality for each prefecture and year, averaged over all ages. In 2011, in Iwate, Miyagi, and Fukushima, there was a large increase in mortality compared to other prefectures. These are northern coastal regions, and the elevated relative mortality is due to the tsunami of 11 March 2011 (see, e.g., Nakahara & Ichikawa, 2013) caused by a magnitude–9.0 earthquake.
In 1995, there was an increase in mortality for Hyogo, which corresponds to the Kobe (Great Hanshin) earthquake (magnitude 6.9) of 17 January 1995. Also evident is the mortality difference in Okinawa up to 2000, suggesting a decline in the relative mortality advantages enjoyed by residents in this region.
The bottom of the panels shows the mortality for each prefecture and age, averaged over all years. Several striking features are apparent. There are differences between prefectures in the care of children and adolescents; this may be due to socio-economic differences and the accessibility of health services (Tanaka et al., 2023; Okamoto et al., 2025).
2.2 Functional cross-correlation plot
Let $d_{t,s}^{g}(u)$ be the age distribution of deaths for age $u$, prefecture $s$, gender $g$ at year $t$, respectively. There are 111 ages, $S=47$ prefectures, and $49$ years. Following Kokoszka et al. (2017) and Mestre et al. (2021), the functional autocorrelation coefficient at lag $h$ can be defined as
$\|C_{h}\|$ $=\int\int C_{h}^{2}(u,u)dudu$
$\rho_{h}$ $=\frac{\|C_{h}\|}{\int C_{0}(u,u)du},$ (1)
where $h$ denotes a lag and $C_{h}(u,u)=\text{cov}[d_{t,s}^{g}(u),d_{t+h,s}^{g}(u)]$. The functional cross-covariance at lag $h$ can be expressed as
$\left\|C_{h}[d_{s}^{g}(u),d_{\text{national}}^{g}(v)]\right\|=\int\int C_{h}^{2}[d_{s}^{g}(u),d_{\text{national}}^{g}(v)]dudv,$
where $C_{h}[d_{s}^{g}(u),d_{\text{national}}^{g}(v)]=\text{cov}[d_{t,s}^{g}(u),d_{t+h,\text{national}}^{g}(v)]$. The functional cross-correlation at lag $h$ is then given by
$\rho_{h}[d_{s}^{g}(u),d^{g}_{\text{national}}(v)]=\frac{\left\|C_{h}[d_{s}^{g}(u),d_{\text{national}}^{g}(v)]\right\|}{\sqrt{\int C_{0}\big[d^{g}_{s}(u),d^{g}_{s}(u)\big]du}\sqrt{\int C_{0}\big[d^{g}_{\text{national}}(v),d^{g}_{\text{national}}(v)\big]dv}}.$ (2)
When $d^{g}_{s}(u)=d_{\text{national}}^{g}(v)$, cross correlation in (2) reduces to autocorrelation in (1).
In Figure 4, we display the functional cross-correlation functions showing the dependence of life-table death counts between each prefecture and the whole country. Each subnational data set shows a stronger cross-correlation with the national data at lags  1 and  2. The cross-correlation reduces as lag increases.
Figure 4: Functional cross-correlation plots showing the dependence of the female, male, and total life-table death counts between each prefecture and the whole country. Prefectures are numbered geographically from north to south.
3 Cumulative distribution function transformation
To model life-table death counts, there exists a range of transformations, including centered log-ratio transformation (Aitchison, 1986), $\alpha$-transformation (Shang & Haberman, 2025c), and CDF transformation (Shang & Haberman, 2025b). Among the three, the CDF transformation can handle not only zero counts but also the monotonicity constraint. In the CDF transformation, we first convert an age distribution of death for each year to a probability by dividing its radix $10^{5}$. Via the cumulative sum, we transform a probability density function into a CDF,
$D_{t,s}^{g}(u_{x})=\sum^{x}_{i=1}d_{t,s}^{g}(u_{i}),\quad x=1,\dots,111,\quad t=1,\dots,n,$
where $D_{t,s}^{g}(u_{111})=1$. In doing so, it enjoys the additional benefit of monotonicity (see also Mayhew & Smith, 2013).
Then, we perform a logistic transformation,
$\mathcal{X}_{t,s}^{g}(u_{x})=\text{logit}[D_{t,s}^{g}(u_{x})]=\ln\Big[\frac{D_{t,s}^{g}(u_{x})}{1-D_{t,s}^{g}(u_{x})}\Big],$
where $\ln(\cdot)$ denotes the natural logarithm. Since $D_{t,s}^{g}(u_{111})=1$, the last column is removed to avoid the undefinedness of the logistic transformation. By considering age as a continuum, the object $\mathcal{X}_{t,s}^{g}(u)$ is an example of a high-dimensional functional time series.
To transform $\mathcal{X}_{t,s}^{g}(u_{x})$ back to the original scale, we first perform an inverse logit transformation, obtaining
$D_{t,s}^{g}(u_{x})=\frac{\exp^{\mathcal{X}_{t,s}^{g}(u_{x})}}{1+\exp^{\mathcal{X}_{t,s}^{g}(u_{x})}},$
where we add back the last column of ones. Then, we take the first-order differencing to obtain
$d_{t,s}^{g}(u_{x})=\Delta_{i=1}^{x}D_{t,s}^{g}(u_{i})\times 10^{5},$
where $\Delta$ represents the first-order differencing, where $\Delta_{i=1}^{1}D_{t,s}^{g}(u_{1})=D_{t,s}^{g}(u_{1})$. The constant $10^{5}$ is the radix of the life-table death count. Since the data are observed discretely, we use cumulative sums and first-order differences to approximate integration and differentiation, respectively.
4 Functional time-series forecasting models
We obtain a set of unconstrained functional time series $(\mathcal{X}_{1,s}^{g}(u),\dots,\mathcal{X}_{n,s}^{g}(u)]$ via the CDF transformation. Generally, we represent it as $\{\mathcal{X}_{t,s}^{g}(u)\}$, which is an arbitrary functional time series defined in a common probability space. It is assumed that the observations $\mathcal{X}_{t,s}^{g}(u)$ are elements of the Hilbert space $\mathcal{H}=\mathcal{L}^{2}(\mathcal{I})$ equipped with the inner product $\langle w,v\rangle=\int_{\mathcal{I}}w(u)v(u)du$, where $u$ represents a continuum and $\mathcal{I}\subset\mathds{R}$ denotes a function support range and $\mathds{R}$ is the real line. Each function is a square-integrable function that satisfies $\|\mathcal{X}_{t,s}^{g}\|^{2}=\int_{\mathcal{I}}(\mathcal{X}_{t,s}^{g}(u))^{2}du<\infty$ and the associated norms.
4.1 Univariate functional time-series (UFTS) forecasting method
We assume $\mathcal{X}_{t,s}^{g}(u)$ has a finite mean and variance. Then, its non-negative definite covariance function is given by
$c_{s}^{g}(u,v)$ $:=\text{Cov}[\mathcal{X}_{s}^{g}(u),\mathcal{X}_{s}^{g}(v)]$
$=\text{E}\left\{[\mathcal{X}_{s}^{g}(u)-\mu_{s}^{g}(u)][\mathcal{X}_{s}^{g}(v)-\mu_{s}^{g}(v)]\right\},\qquad\forall u,v\in\mathcal{I},$
where $\mu_{s}^{g}(u)=\text{E}\{\mathcal{X}_{t,s}^{g}(u)\}$. By Mercer’s lemma, there exists a set of orthonormal sequences $(\phi_{k})$ of continuous functions in $\mathcal{L}^{2}(\mathcal{I})$ and a non-increasing sequence $\lambda_{k,s}^{g}$ of positive numbers such that
$c_{s}^{g}(u,v)=\sum^{\infty}_{k=1}\lambda_{k,s}^{g}\phi_{k,s}^{g}(u)\phi_{k,s}^{g}(v),$
where $\lambda_{k,s}^{g}$ and $\phi_{k,s}^{g}(u)$ denote the $k$th eigenvalue and eigenfunction, respectively. The Karhunen-Loève expansion of a stochastic process $\mathcal{X}_{t,s}^{g}(u)$ can be expressed as
$\mathcal{X}_{t,s}^{g}(u)$ $=\mu_{s}^{g}(u)+\sum_{k=1}^{\infty}\beta_{t,k,s}^{g}\phi_{k,s}^{g}(u)$
$=\mu_{s}^{g}(u)+\sum_{k=1}^{K_{s}^{g}}\beta_{t,k,s}^{g}\phi_{k,s}^{g}(u)+e_{t,s}^{g}(u),$
where the principal component score $\beta_{t,k,s}^{g}$ is given by the projection of $\mathcal{X}_{t,s}^{g}(u)-\mu_{s}^{g}(u)]$ in the direction of the $k$th eigenfunction $\phi_{k,s}^{g}(u)$, i.e., $\beta_{t,k,s}^{g}=\langle\mathcal{X}_{t,s}^{g}(u)-\mu_{s}^{g}(u),\phi_{k,s}^{g}(u)\rangle$; and $e_{t,s}^{g}(u)$ denotes the model error function with a mean of zero and a finite variance, and $K_{s}^{g}<n$ is the number of retained components. The eigenfunctions represent the fixed shapes of variation (e.g., longevity shift), while the scores capture the time-varying magnitude of these shifts.
There are a number of ways to select the retained number of principal components, such as the bootstrap approach proposed by Hall & Vial (2006) and Bathia et al. (2010), description length approach proposed by Poskitt & Sengarapillai (2013), pseudo-Akaike information criterion (Shibata, 1981), scree plot (Cattell, 1966), and eigenvector variability plot (Tu et al., 2009). We consider the eigenvalue ratio criterion (EVR) of Li et al. (2020), defined as
$K_{s}^{g}=\operatorname*{arg\,min}_{k=1,\dots,k_{\max}}\left\{\frac{\widehat{\lambda}_{k+1,s}^{g}}{\widehat{\lambda}_{k,s}^{g}}\times\mathds{1}\Big\{\frac{\widehat{\lambda}_{k,s}^{g}}{\widehat{\lambda}_{1,s}^{g}}\geq\eta\Big\}+\mathds{1}\Big\{\frac{\widehat{\lambda}_{k,s}^{g}}{\widehat{\lambda}_{1,s}^{g}}<\eta\Big\}\right\},$ (3)
where $\mathds{1}\{\cdot\}$ is the binary indicator function, $\eta=\frac{1}{\ln[\max(\widehat{\lambda}_{1,s}^{g},n)]}$ is a small positive value. Instead of searching through $n$, we restrict our searching range by setting $k_{\max}=\#\{k|\widehat{\lambda}_{k,s}^{g}\geq\frac{1}{n}\sum^{n}_{k=1}\widehat{\lambda}_{k,s}^{g},k\geq 1\}$. For comparison, we also consider $K_{s}^{g}=6$ as suggested in Hyndman et al. (2013). Note that when $K_{s}^{g}=1$, it reduces to the Lee & Carter’s (1992) model.
Conditioning on the observed functions $\bm{\mathcal{X}}_{s}^{g}(u)=[\mathcal{X}_{1,s}^{g}(u),\dots,\mathcal{X}_{n,s}^{g}(u)]$ and the estimated functional principal components $\bm{B}=\big\{\phi_{1,s}^{g}(u),\dots,\phi_{K,s}^{g}(u)\big\}$, the $h$-step-ahead point forecast of $\mathcal{X}_{n+h,s}^{g}(u)$ can be expressed as
$\widehat{\mathcal{X}}_{n+h|n,s}^{g}(u)=\text{E}\big[\mathcal{X}_{n+h,s}^{g}(u)|\bm{\mathcal{X}}_{s}^{g}(u),\bm{B}\big]=\widehat{\mu}_{s}^{g}(u)+\sum^{K_{s}^{g}}_{k=1}\widehat{\beta}_{n+h|n,k,s}^{g}\widehat{\phi}_{k,s}^{g}(u),$
where $\widehat{\mu}_{s}^{g}(u)=\frac{1}{n}\sum^{n}_{t=1}\mathcal{X}_{t,s}^{g}(u)$ and $\widehat{\beta}_{n+h|n,k,s}^{g}$ denotes time-series forecasts of the $k$th principal component scores. The forecasts of these scores can be obtained via a univariate time-series forecasting method, such as exponential smoothing (Hyndman et al., 2008). Computationally, we implement the ets function in the forecast package (Hyndman & Khandakar, 2008), which selects the optimal parameters based on the corrected Akaike information criterion.
4.2 Multivariate functional time-series (MFTS) forecasting method
The univariate functional time-series method does not account for correlations between female and male data within the same region. However, explicitly modeling the correlation between multiple series may improve forecast accuracy. We consider a multivariate functional time-series method to jointly model and forecast multiple series that may be correlated.
Let $\mathcal{X}_{t,s}^{F}(u)$ and $\mathcal{X}_{t,s}^{M}(u)$ represent female and male subnational age distribution of death counts. As our multiple functional time series have the same function support, we consider data where each observation consists of $\omega\geq 2$ functions, that is, $\mathcal{X}_{t,s}^{F}(u),\mathcal{X}_{t,s}^{M}(u)]\in\mathds{R}^{2}$, where $u\in\mathcal{I}$.
Multivariate functional time series are stacked into a vector. Let $\nu^{F}_{s}(u)$ and $\nu^{M}_{s}(u)$ represent the mean function, respectively. For $u,v\in\mathcal{I}$, the theoretical cross-covariance function can be defined with elements
$c_{s}(u,v):=\text{Cov}[\mathcal{X}^{F}_{s}(u),\mathcal{X}^{M}_{s}(v)]=\text{E}\left\{[\mathcal{X}^{F}_{s}(u)-\nu^{F}_{s}(u)][\mathcal{X}^{M}_{s}(u)-\nu^{M}_{s}(u)]\right\}.$
The Karhunen-Loève expansion of a stochastic process can be expressed as
$\bm{\mathcal{X}}_{s}(u)=\bm{\nu}_{s}(u)+\bm{\Phi}_{s}(u)\bm{\beta}_{t,s}^{\top},$
where $\bm{\mathcal{X}}_{s}(u)=[\bm{\mathcal{X}}_{s}^{F}(u),\bm{\mathcal{X}}_{s}^{M}(u)]$, $\bm{\mathcal{X}}_{s}^{F}(u)=[\mathcal{X}_{1,s}^{F}(u),\dots,\mathcal{X}_{n,s}^{F}(u)]$ denote stacked historical functions, and $\bm{\nu}_{s}(u)=[\bm{\nu}_{s}^{F}(u),\bm{\nu}_{s}^{M}(u)]$. Moreover, $\bm{\beta}_{s}=(\bm{\beta}_{1,s},\dots,\bm{\beta}_{K,s})$ and $\bm{\beta}_{1,s}=(\beta_{1,1,s},\dots,\beta_{n,1,s})$ are the vectors of the principal component scores, and
$\bm{\Phi}_{s}(u)=\left[\begin{array}[]{cccccc}\phi_{1,s}^{F}(u)&\cdots&\phi_{K,s}^{F}(u)&&\cdots&\\ &\cdots&&\phi_{1,s}^{M}(u)&\cdots&\phi_{K,s}^{M}(u)\end{array}\right]_{2\times(K\times 2)},$
where $K$ denotes the retained number of principal components shared by females and males.
Conditioning on the previous functions $\bm{\mathcal{X}}_{s}(u)$ and the estimated functional principal components $\bm{\Phi}_{s}(u)$, the $h$-step-ahead point forecast of $\bm{\mathcal{X}}_{n+h,s}(u)$ can be expressed as
$\widehat{\bm{\mathcal{X}}}_{n+h|n,s}(u)=\text{E}\left[\bm{\mathcal{X}}_{n+h,s}(u)|\bm{\mathcal{X}}_{s}(u),\bm{\Phi}_{s}(u)\right]=\widehat{\nu}_{s}(u)+\widehat{\bm{\Phi}}_{s}(u)\widehat{\bm{\beta}}_{n+h|n,s}^{\top},$
where $\widehat{\bm{\beta}}_{n+h|n,s}$ denotes the time-series forecasts of the principal component scores corresponding to the female and male series, $\widehat{\nu}_{s}(u)$ and $\widehat{\bm{\Phi}}_{s}(u)$ denote the estimated mean function and estimated functional principal components, respectively.
4.3 Multilevel functional time-series (MLFTS) forecasting method
The multilevel functional data model strongly resembles the two-way functional analysis of variance studied by Morris & Carroll (2006) and Cuesta-Albertos & Febrero-Bande (2010). The basic idea is to extract common patterns shared by multiple series $R_{t,s}(u)$ and series-specific patterns $U_{t,s}^{g}(u)$. The common trend and series-specific trend are modeled by projecting them onto the eigenvectors of the covariance functions of the aggregated and series-specific stochastic processes, respectively. For $t=1,\dots,n$, a curve can be expressed as
$\mathcal{X}_{t,s}^{g}(u)=\mu^{g}_{s}(u)+R^{c}_{t,s}(u)+U_{t,s}^{g}(u),$ (4)
where $R^{c}_{t,s}(u)$ can be the simple average of the female and male series. To ensure model and parameter identifiability, we implement a two-stage estimation procedure based on functional principal component decomposition.
Since the stochastic processes $R^{c}_{t,s}(u)$ and $U^{g}_{t,s}(u)$ are unknown in practice, the population eigenvalues and eigenfunctions can only be approximated at best through a set of realizations $\bm{R}^{c}_{s}(u)$ and $\bm{U}_{s}^{g}(u)$. From the covariance function of $\bm{R}^{c}_{s}(u)$, we can extract a set of functional principal components and their associated scores, along with a set of residual functions. From the covariance function of the residuals, we can then extract a second set of functional principal components and their associated scores. While the first functional principal component decomposition captures the common pattern across multiple series, the second functional principal component decomposition captures the series-specific residual trend.
The sample versions of the series-specific mean function, a common trend, and the series-specific residual trend for a set of functional time series can be estimated by
$\widehat{\mu}_{s}^{g}(u)$ $=\frac{1}{n}\sum^{n}_{t=1}\mathcal{X}_{t,s}^{g}(u)$ (5)
$R_{t,s}^{c}(u)$ $\approx\sum^{K}_{k=1}\beta_{t,k,s}^{c}\phi_{k,s}^{c}(u)$ (6)
$U_{t,s}^{g}(u)$ $\approx\sum^{L}_{l=1}\gamma_{t,l,s}^{g}\psi_{l,s}^{g}(u),$ (7)
where $\widehat{\mu}^{g}_{s}(u)$ represents the simple average of the female or male series, $\bm{\beta}_{k,s}^{c}=(\beta_{1,k,s}^{c},\dots,\beta_{n,k,s}^{c})$ represents the $k$th sample principal component scores of $\bm{R}^{c}_{s}(u)$, and $\bm{\gamma}_{l,s}^{g}=(\gamma_{1,l,s}^{g},\dots,\gamma_{n,l,s}^{g})$ represents the $l$th sample principal component scores of $\bm{U}_{s}^{g}(u)$.
Plugging (5) into (7) into (4), we obtain
$\mathcal{X}_{t,s}^{g}(u)\approx\widehat{\mu}^{g}_{s}(u)+\sum^{K}_{k=1}\beta_{t,k,s}^{c}\phi_{k,s}^{c}(u)+\sum^{L}_{l=1}\gamma_{t,l,s}^{g}\psi_{l,s}^{g}(u)+e_{t,s}^{g}(u),$
where $e_{t,s}^{g}(u)$ denotes measurement error with a finite variance.
To select the retained number of components, we use the eigenvalue ratio criterion in (3). A feature of the multilevel method is its ability to estimate the proportion of variability explained by aggregated data. For a given population $g$, a measure of the within-cluster variability is given by
$\frac{\sum_{k=1}^{K}\widehat{\lambda}_{k,s}^{c}}{\sum_{k=1}^{K}\widehat{\lambda}_{k,s}^{c}+\sum_{l=1}^{L}\widehat{\lambda}_{l,s}^{g}}.$
When the common trend explains the primary mode of total variability, the within-cluster variability is close to 1.
[…]
$\widehat{\mathcal{X}}_{n+h|n,s}^{g}(u)$ $=\text{E}[\mathcal{X}_{n+h,s}^{g}(u)|\mu^{g}_{s}(u),\bm{\mathcal{X}}^{g}_{s}(u),\bm{\Phi}_{s}(u),\bm{\Psi}_{s}^{g}(u)]$
$=\widehat{\mu}^{g}_{s}(u)+\sum^{K}_{k=1}\widehat{\beta}_{n+h|n,k,s}^{c}\widehat{\phi}_{k,s}^{c}(u)+\sum^{L}_{l=1}\widehat{\gamma}_{n+h|n,l,s}^{g}\widehat{\psi}_{l,s}^{g}(u),$
where $\widehat{\mu}_{s}^{g}(u)=\frac{1}{n}\sum^{n}_{t=1}\mathcal{X}_{t,s}^{g}(u)$, $\{\widehat{\beta}_{n+h|n,k,s}^{c},k=1,\dots,K\}$ and $\{\widehat{\gamma}_{n+h|n,l,s}^{g},l=1,\dots,L\}$ are the forecasted principal component scores, obtained from a univariate time-series forecasting method.
4.4 Two-way functional analysis of variance (FANOVA
To model $\mathcal{X}_{t,s}^{g}(u)$, we consider a two-way FANOVA decomposition, expressed as:
$\mathcal{X}_{t,s}^{g}(u)=\mu(u)+\alpha_{s}(u)+\beta^{g}(u)+\epsilon_{t,s}^{g}(u),$
where $\mu(u)$ denotes the functional grand effect, $\alpha_{s}(u)$ denotes the functional row effect, $\beta^{g}(u)$ denotes the functional column effect, and $\epsilon_{t,s}^{g}(u)$ denotes the error function.
The FANOVA model can be estimated by means with
$\widehat{\mu}(u)$ $=\frac{1}{n_{s}\times 2\times n}\sum^{n_{s}}_{s=1}\sum^{2}_{g=1}\sum^{n}_{t=1}\mathcal{X}_{t,s}^{g}(u)$
$\widehat{\alpha}_{s}(u)$ $=\frac{1}{2\times n}\sum^{2}_{g=1}\sum^{n}_{t=1}\mathcal{X}_{t,s}^{g}(u)-\widehat{\mu}(u)$
$\widehat{\beta}^{g}(u)$ $=\frac{1}{n_{s}\times n}\sum^{n_{s}}_{s=1}\sum^{n}_{t=1}\mathcal{X}_{t,s}^{g}(u)-\widehat{\mu}(u).$
Some identifiability constraints exist, $\forall u\in\mathcal{I}$, $\sum^{n_{s}}_{s=1}\alpha_{s}(u)=\sum^{2}_{g=1}\beta^{g}(u)=0$, and $\sum^{n_{s}}_{s=1}\mathcal{X}_{t,s}^{g}(u)=\sum^{2}_{g=1}\mathcal{X}_{t,s}^{g}(u)=0$, $\forall t$. The FANOVA model decomposes the functional time series (FTS) jointly for all prefectures and genders into three parameters: the overall mean $\widehat{\mu}(u)$, which represents the grand effect measuring the overall baseline age distribution of deaths; $\widehat{\alpha}_{s}(u)$, which represents the functional prefecture effect measuring the time-invariant functional difference attribute to each prefecture, relative to the grand mean; and $\widehat{\beta}^{g}(u)$, which represents the functional gender effect measuring the time-invariant functional difference between genders, relative to the grand mean.
The residual functions $\widehat{\epsilon}_{t,s}^{g}(u)=\mathbf{X}_{t,s}^{g}(u)-\widehat{\mu}(u)-\widehat{\alpha}_{s}(u)-\widehat{\beta}^{g}(u)$ are time-varying, for $t=1,\dots,n$. Following the methodology of Jiménez-Varón et al. (2024), we model and forecast these residual functions using the multivariate functional time series (MFTS) approach.
4.5 Two-stage functional principal component analyses
Gao et al. (2019) propose a two-stage functional principal component decomposition to model and forecast HDFTS. One of the earliest methods for comparison (see, e.g., Tavakoli et al., 2023; Shang, 2025), it consists of three parts.
  1. 1)
    Functional principal component analysis is performed separately on each set of functional time series, resulting in $P$ sets of principal component scores of low dimension $p_{0}$.
  2. 2)
    The first functional principal component scores from each of $P$ sets of functional time series are combined into an $P\times 1$ vector. We fit factor models to the scores to further reduce the dimension into an $r\times 1$ vector, where $r<<P$. The same is carried out for the second, third, and so on until the $p_{0}$th scores. The vector of the $P$ functional time series is reduced to an $r\times p_{0}$ matrix consisting of factors.
  3. 3)
    A univariate time series model can be fitted to each factor, and forecasts are generated. Forecast factors can be used to construct forecast functions.
5 Construction of prediction intervals
We split the data sample consisting of 49 years from 1975 to 2023 into a training set consisting of 16 years, a validation set consisting of 16 years, and a testing set consisting of 17 years. While the choice of different data splits is arbitrary, it is customary in forecasting to allocate at least one-third of the total data for evaluation (Hyndman & Athanasopoulos, 2021).
Using the data in the training sample, we implement an expanding-window forecast scheme to obtain $h$-step-ahead density forecasts in the validation set for $h=1,2,\dots,16$. For each horizon, we have a different number of curves in the validation set. For instance, when $h=1$, we have 16 years to evaluate the forecast errors, denoted by $\widehat{\epsilon}_{m,s}^{g}(u)=d_{m,s}^{g}(u)-\widehat{d}_{m,s}^{g}(u)$ for $m=1,2,\dots,M=16$; when $h=16$, we have two years to evaluate the residual functions between the samples in the validation set and their corresponding forecasts. From these residual functions, we compute the functional standard deviation, denoted by
$\gamma(u)=\text{sd}[\widehat{\epsilon}_{1,s}^{g}(u),\dots,\widehat{\epsilon}_{M,s}^{g}(u)].$ (8)
We aim to determine $(\overline{\xi}_{\alpha},\underline{\xi}_{\alpha})$ such that $\alpha\times 100\%$ of the residuals satisfy
$\underline{\xi}_{\alpha}\gamma(u)\leq\widehat{\epsilon}_{m,s}^{g}(u)\leq\overline{\xi}_{\alpha}\gamma(u),$
where $(\overline{\xi}_{\alpha},\underline{\xi}_{\alpha})$ are the tuning parameters.
Typically, the constants $\overline{\xi}_{\alpha}$ and $\underline{\xi}_{\alpha}$ are chosen equal. By the law of large numbers, one may achieve
$\text{Pr}\left[-\xi_{\alpha}\gamma(u)\leq d_{n+h,s}^{g}(u)-\widehat{d}_{n+h,s}^{g}(u)\leq\xi_{\alpha}\gamma(u)\right]\approx\frac{1}{M}\sum^{M}_{m=1}\mathds{1}\left[-\xi_{\alpha}\gamma(u)\leq\widehat{\epsilon}_{m,s}^{g}(u)\leq\xi_{\alpha}\gamma(u)\right],$
where $M$ denotes the number of years in the validation set.
To determine the optimal $\xi_{\alpha}$, the samples in the validation set are used to calibrate the prediction intervals so that the empirical coverage probabilities defined in (10) are close to their nominal coverage probabilities.
In Appendix C, we also present the conformal prediction interval approach of Shafer & Vovk (2008), which does not require a tuning parameter. By considering a quantile of the absolute residual functions, the pointwise prediction interval for the out-of-sample testing set is centered on the point forecast, with its spread determined by the chosen quantile.
6 Results
6.1 Expanding window scheme
An expanding window analysis of a time-series model is commonly used to assess model and parameter stability and prediction accuracy over time. The expanding window analysis assesses the stability of a model’s parameters by computing parameter estimates and their forecasts over an expanding window across the sample (for details, see Zivot & Wang, 2006, pp. 313–314).
Using the first 32 years from 1975 to 2006 in the Japanese subnational age- and sex-specific life-table death counts, we produce one- to 16-step-ahead forecasts. Through an expanding-window approach, we re-estimate the parameters in the time-series forecasting models using the first 33 years from 1975 to 2007. Forecasts from the estimated models are then produced for one- to 16-step-ahead forecasts. We iterate this process by increasing the sample size by one year until reaching the end of the data period in 2023. This process produces 16 one-step-ahead forecasts, 15 two-step-ahead forecasts, $\dots$, and one 16-step-ahead forecast. We compare these forecasts with holdout samples to assess out-of-sample forecast accuracy. In Figure 5, we display a diagram of the expanding window forecast scheme for the forecast horizon $h=1$, although we also consider other forecast horizons from $h=2$ to $16$.
[…]
Figure 5: A diagram of the expanding-window forecast scheme.
6.2 Point forecast error metrics
Since life-table death counts can be considered probability density functions, we use some density-evaluation measures. These measures include the discrete version of the KLD (Kullback & Leibler, 1951), the square root of the JSD (Shannon, 1948). For two probability density functions, denoted by $d_{m+\xi}$ and $\widehat{d}_{m+\xi}$, the symmetric version of the KLD, also known as Jeffrey divergence, is defined as
$\text{KLD}(h)=\$$\text{D}_{\text{KL}}(d_{m+\xi}||\widehat{d}_{m+\xi})+\text{D}_{\text{KL}}(\widehat{d}_{m+\xi}||d_{m+\xi})$
$=\$$\frac{1}{111\times(17-h)}\sum^{16}_{\xi=h}\sum_{x=1}^{111}d_{m+\xi}(u_{x})\cdot\left[\ln d_{m+\xi}(u_{x})-\ln\widehat{d}_{m+\xi}(u_{x})\right]+$
$\frac{1}{111\times(17-h)}\sum^{16}_{\xi=h}\sum_{x=1}^{111}\widehat{d}_{m+\xi}(u_{x})\cdot\left[\ln\widehat{d}_{m+\xi}(u_{x})-\ln d_{m+\xi}(u_{x})\right],$ (9)
which is non-negative, for $h=1,2,\dots,16$.
An alternative metric is the JSD, defined by
$\text{JSD}(h)=\frac{1}{2}D_{\text{KL}}(d_{m+\xi}||\delta_{m+\xi})+\frac{1}{2}D_{\text{KL}}(\widehat{d}_{m+\xi}||\delta_{m+\xi}),$
where $\delta_{m+\xi}$ measures a common quantity between $d_{m+\xi}$ and $\widehat{d}_{m+\xi}$. Following the early work of Shang & Haberman (2020), we consider the geometric mean $\delta_{m+\xi}=\sqrt{d_{m+\xi}\widehat{d}_{m+\xi}}$. To make the JSD a metric between two probability densities, we take the square root of the JSD (see, e.g., Fuglede & Topsoe, 2004).
6.3 Comparison of point forecast accuracy
In Table LABEL:tab:1, we compare the point forecast accuracy across the functional time-series models, averaged across 47 prefectures. Based on the summary statistics of the KLD and JSD, the HDFPCA model provides the most accurate point forecasts for the female data. This result indicates that the two-stage principal component analyses can adequately summarize the main features in the female data without much loss of information. The female data are relatively stable in comparison to the male counterpart. For the male data, the MFTS and MLFTS models provide the most accurate point forecasts. For the UFTS, MFTS, MLFTS, and FANOVA models, the number of retained functional principal components was determined by the EVR criterion. For the HDFPCA model, we follow the default setting of Gao et al. (2019), where the number of retained components is six and two in the first and second stages, respectively. In Appendix A, we also provide the results for $K=6$ as a sensitivity analysis.
Table 1: Averaged across 47 prefectures, we evaluate and compare the point forecast accuracy of the functional time-series models, measured by the KLD and JSD. The number of components is determined by the EVR criterion.
Female Male
Metric $h$ UFTS MFTS MLFTS FANOVA HDFPCA UFTS MFTS MLFTS FANOVA HDFPCA
KLD 1 0.004 0.010 0.004 0.006 0.007 0.004 0.004 0.003 0.005 0.003
2 0.007 0.013 0.006 0.007 0.007 0.005 0.006 0.005 0.005 0.003
3 0.008 0.015 0.007 0.009 0.009 0.006 0.007 0.005 0.006 0.004
4 0.010 0.017 0.009 0.010 0.006 0.006 0.007 0.006 0.007 0.005
5 0.013 0.019 0.010 0.012 0.013 0.007 0.008 0.007 0.007 0.006
6 0.015 0.021 0.011 0.013 0.007 0.008 0.009 0.007 0.008 0.007
7 0.018 0.023 0.012 0.015 0.019 0.009 0.010 0.008 0.008 0.009
8 0.022 0.025 0.014 0.018 0.008 0.010 0.011 0.009 0.009 0.012
9 0.027 0.027 0.017 0.022 0.025 0.011 0.012 0.010 0.010 0.015
10 0.033 0.030 0.021 0.026 0.010 0.012 0.014 0.011 0.011 0.018
11 0.040 0.034 0.024 0.033 0.032 0.012 0.015 0.012 0.012 0.022
12 0.048 0.039 0.030 0.040 0.014 0.013 0.017 0.014 0.013 0.026
13 0.059 0.047 0.036 0.047 0.039 0.014 0.020 0.015 0.014 0.030
14 0.068 0.040 0.042 0.061 0.019 0.014 0.012 0.015 0.017 0.035
15 0.089 0.054 0.060 0.080 0.044 0.015 0.015 0.021 0.022 0.039
16 0.118 0.070 0.079 0.106 0.019 0.015 0.019 0.029 0.030 0.040
17 0.154 0.084 0.093 0.120 0.038 0.015 0.020 0.034 0.034 0.039
Mean 0.043 0.033 0.028 0.037 0.019 0.010 0.012 0.012 0.013 0.019
Median 0.027 0.027 0.017 0.022 0.014 0.011 0.012 0.010 0.010 0.015
JSD 1 0.034 0.049 0.034 0.039 0.042 0.030 0.032 0.028 0.034 0.029
2 0.041 0.055 0.040 0.043 0.038 0.035 0.036 0.033 0.036 0.030
3 0.045 0.059 0.043 0.046 0.050 0.037 0.039 0.036 0.038 0.032
4 0.049 0.063 0.046 0.049 0.037 0.040 0.041 0.038 0.040 0.035
5 0.054 0.066 0.049 0.053 0.060 0.042 0.043 0.040 0.041 0.039
6 0.058 0.069 0.052 0.056 0.040 0.044 0.045 0.041 0.043 0.043
7 0.063 0.071 0.055 0.060 0.070 0.047 0.047 0.043 0.044 0.049
8 0.069 0.073 0.059 0.065 0.044 0.050 0.049 0.045 0.046 0.055
9 0.076 0.075 0.063 0.070 0.082 0.052 0.051 0.048 0.047 0.062
10 0.083 0.079 0.069 0.076 0.051 0.054 0.053 0.050 0.049 0.069
11 0.092 0.083 0.075 0.085 0.092 0.056 0.054 0.052 0.052 0.075
12 0.102 0.088 0.082 0.096 0.058 0.058 0.056 0.055 0.054 0.081
13 0.116 0.094 0.092 0.106 0.100 0.060 0.059 0.058 0.057 0.087
14 0.128 0.097 0.102 0.121 0.066 0.061 0.057 0.061 0.064 0.094
15 0.150 0.112 0.122 0.141 0.104 0.062 0.062 0.070 0.072 0.097
16 0.176 0.126 0.141 0.164 0.068 0.064 0.067 0.081 0.084 0.095
17 0.202 0.134 0.152 0.175 0.094 0.064 0.069 0.087 0.088 0.093
Mean 0.090 0.082 0.075 0.085 0.064 0.050 0.051 0.051 0.052 0.063
Median 0.076 0.075 0.063 0.070 0.060 0.052 0.051 0.048 0.047 0.062
Figure 6 presents two heatmaps comparing seven functional time-series models across 16 forecast horizons for both female and male data. For each horizon, we count how often each method produces the smallest point forecast errors, as measured by the KLD. Choosing $K=6$ has a slight advantage compared to the one based on the EVR criterion. For shorter forecast horizons, the UFTS, MFTS, and MLFTS models perform equally well for the female data. However, for relatively longer forecast horizons, the HDFPCA model is recommended. For male data, the MLFTS model is generally preferred; only at the shorter forecast horizon is $K=6$ advantageous. For relatively longer forecast horizons, the FANOVA with the EVR criterion is recommended.
(a) Female data
(b) Male data
Figure 6: Heatmaps comparing the point forecast accuracy, measured by the KLD, across 17 forecast horizons for nine functional time-series models on the testing female and male data. Since there are 47 prefectures, each row sums to 47. A method that appears more than 5 times is highlighted in green, while those that appear fewer than 5 times are shown in red. If a method appears exactly five times, it is displayed in white.
To formally assess whether the observed differences in point forecast accuracy obtained via the different presented methods are statistically significant, we further apply the Model Confidence Set (MCS) procedure (Hansen et al., 2011) based on the Kullback–Leibler divergence across all forecast horizons and prefectures. The detailed MCS results and corresponding heatmap visualizations are reported in Appendix D.
6.4 Interval forecast error metrics
To evaluate interval forecast accuracy, we consider the CPD between the empirical coverage probability (ECP) and nominal coverage probability, as well as the mean interval score of Gneiting & Raftery (2007). For each year in the forecasting period, the $h$-step-ahead prediction intervals are computed at the $100(1-\alpha)\%$ nominal coverage probability. We consider the common case of the symmetric $100(1-\alpha)\%$ prediction intervals, with lower and upper bounds that are quantiles at $\alpha/2$ and $1-\alpha/2$, denoted by $\widehat{d}_{n+\xi,s}^{g,\text{lb}}(u)$ and $\widehat{d}_{n+\xi,s}^{g,\text{ub}}(u)$. The ECP and CPD are defined as
$\text{ECP}_{h}$ $=\frac{1}{111\times(17-h)}\times\sum_{\xi=h}^{16}\sum_{u=1}^{111}\mathds{1}\left\{\widehat{d}_{\nu+\xi,s}^{g,\text{lb}}(u)\leq d_{\nu+\xi,s}^{g}(u)\leq\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u)\right\},$ (10)
$\text{CPD}_{h}$ $=\frac{1}{111\times(17-h)}\times\sum_{\xi=h}^{16}\sum_{u=1}^{111}\left[\mathds{1}\{d_{\nu+\xi,s}^{g}(u)>\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u)\}+\mathds{1}\{d_{\nu+\xi,s}^{g}(u)<\widehat{d}_{\nu+\xi,s}^{g,\text{lb}}(u)\}\right],$
where $h=1,\dots,16$, $\nu$ denotes the end of the training set when we calibrate the CPD or interval score, and $\nu$ denotes the end of the validation set when we evaluate the interval forecast accuracy.
For different ages and years in the testing set, the mean and median CPD are defined as
$\overline{\text{CPD}}$ $=\frac{1}{16}\sum^{16}_{h=1}\text{CPD}_{h},$
M[CPD] $=\text{median}(\text{CPD}_{h}).$
As defined by Gneiting & Raftery (2007), a scoring rule for the prediction intervals at age $u$ is
$S_{\alpha,\xi}\left[\widehat{d}^{g,\text{lb}}_{\nu+\xi,s}(u),\widehat{d}^{g,\text{ub}}_{\nu+\xi,s}(u),d^{g}_{\nu+\xi,s}(u)\right]$ $=\left[\widehat{d}^{g,\text{ub}}_{\nu+\xi,s}(u)-\widehat{d}^{g,\text{lb}}_{\nu+\xi,s}(u)\right]$
$+\frac{2}{\alpha}\left[\widehat{d}^{g,\text{lb}}_{\nu+\xi,s}(u)-d_{n+\xi,s}^{g}(u)\right]\mathds{1}\left\{d_{\nu+\xi,s}^{g}(u)<\widehat{d}_{\nu+\xi,s}^{g,\text{lb}}(u)\right\}$
$+\frac{2}{\alpha}\left[d_{\nu+\xi,s}^{g}(u)-\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u)\right]\mathds{1}\left\{d_{\nu+\xi,s}^{g}(u)>\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u)\right\},$
where the level of significance is customarily set to be $\alpha=0.2$ or $0.05$. The interval score rewards a narrow prediction interval width if and only if $100(1-\alpha)\%$ of the holdout densities lie within the prediction interval.
For different ages and years in the testing set, the mean interval score is defined by
$\overline{S}_{\alpha}(h)=\frac{1}{111\times(17-h)}\times\sum^{16}_{\xi=h}\sum^{111}_{u=1}S_{\alpha,\xi}\left[\widehat{d}_{\nu+\xi,s}^{g,\text{lb}}(u),\widehat{d}_{\nu+\xi,s}^{g,\text{ub}}(u),d_{\nu+\xi,s}^{g}(u)\right].$
Averaging over all forecast horizons, we obtain the overall mean interval score
$\overline{S}_{\alpha}=\frac{1}{16}\sum^{16}_{h=1}\overline{S}_{\alpha}(h),\qquad\text{M}[S_{\alpha}]=\text{median}[\overline{S}_{\alpha}(h)].$
6.5 Comparison of interval forecast accuracy
In Table LABEL:tab:2, averaged across 47 prefectures, we compare the interval forecast accuracy among the functional time-series models using the EVR criterion. From the summary statistics of the CPD and interval score at the level of significance $\alpha=0.2$, the MFTS offers the smallest CPD value for the female data, while the UFTS and MLFTS offer the smallest CPD values for the male data. In terms of the interval scores, the UFTS and MFTS provide the smallest interval scores for the female data, while the FANOVA gives the smallest interval scores for the male data. In Appendix B, we also provide the results for $K=6$.
Table 2: Averaged across 47 prefectures, we evaluate and compare the interval forecast accuracy, as measured by the CPD and interval score, for the seven functional time-series models at the significance level $\alpha=0.2$. The number of components is determined by the EVR criterion.
Female Male
Metric $h$ UFTS MFTS MLFTS FANOVA HDFPCA UFTS MFTS MLFTS FANOVA HDFPCA
$\overline{\text{CPD}}_{0.2}$ 1 0.061 0.080 0.047 0.075 0.077 0.060 0.100 0.043 0.079 0.067
2 0.060 0.074 0.046 0.085 0.054 0.060 0.102 0.047 0.090 0.084
3 0.058 0.068 0.050 0.096 0.091 0.058 0.105 0.055 0.094 0.086
4 0.058 0.064 0.055 0.110 0.067 0.058 0.108 0.056 0.097 0.087
5 0.053 0.056 0.057 0.113 0.094 0.053 0.101 0.056 0.096 0.085
6 0.060 0.050 0.065 0.124 0.085 0.060 0.104 0.057 0.097 0.092
7 0.061 0.048 0.069 0.125 0.084 0.061 0.097 0.057 0.091 0.094
8 0.061 0.046 0.074 0.126 0.095 0.061 0.089 0.055 0.085 0.087
9 0.068 0.047 0.081 0.128 0.072 0.068 0.083 0.062 0.076 0.084
10 0.071 0.047 0.086 0.139 0.123 0.071 0.085 0.069 0.084 0.090
11 0.078 0.049 0.093 0.145 0.059 0.078 0.087 0.075 0.085 0.084
12 0.075 0.053 0.092 0.146 0.126 0.075 0.083 0.078 0.085 0.089
13 0.076 0.054 0.084 0.147 0.065 0.076 0.073 0.077 0.089 0.081
14 0.073 0.068 0.087 0.147 0.100 0.073 0.078 0.084 0.095 0.099
15 0.069 0.090 0.091 0.163 0.098 0.069 0.080 0.088 0.087 0.105
16 0.062 0.098 0.093 0.178 0.062 0.062 0.077 0.081 0.085 0.088
Mean 0.065 0.062 0.073 0.128 0.085 0.065 0.091 0.065 0.089 0.087
Median 0.062 0.055 0.078 0.127 0.084 0.062 0.088 0.060 0.088 0.086
$\overline{S}_{0.2}$ 1 268 256 433 287 325 256 279 218 258 248
2 295 302 466 336 335 295 306 243 279 281
3 317 347 494 389 402 317 332 265 293 306
4 345 397 524 450 357 345 359 285 312 327
5 368 445 551 508 457 368 379 300 321 344
6 398 510 591 587 394 398 401 319 338 370
7 420 583 618 672 547 420 415 339 354 394
8 443 675 651 780 450 443 432 361 371 420
9 465 763 688 885 633 465 446 386 386 441
10 488 875 750 1033 492 488 473 409 398 484
11 500 998 761 1191 775 500 491 439 421 513
12 520 1159 862 1377 542 520 518 461 448 585
13 527 1306 914 1572 1020 527 470 482 473 658
14 563 1533 1058 1877 617 563 521 571 535 783
15 627 1869 1225 2203 1432 627 588 700 621 924
16 843 2151 1456 2526 848 843 807 991 865 1139
Mean 481 481 784 1281 705 481 491 453 440 534
Median 443 447 688 885 471 443 446 439 421 431
In Figure 7, we present two heatmaps comparing seven functional time-series models across 16 forecast horizons for both female and male testing data. For each horizon, we count how often each method produces the smallest interval forecast errors, as measured by the CPD and interval score. Based on the CPD results, we recommend the MFTS for the female data and the MLFTS for the male data, especially for $K=6$. By examining the interval score results, we recommend the MFTS for shorter horizons and the HDFPCA for longer horizons in the female data. For the male data, the MFTS generally provides the smallest interval scores.
(a) CPD for females
(b) CPD for males
(c) Interval score for females
(d) Interval score for males
Figure 7: Heatmaps comparing the interval forecast accuracy at the significance level $\alpha=0.2$, measured by the CPD and interval score, across seven functional time-series models and 16 forecast horizons for the female and male testing data. Since there are 47 prefectures, each row sums to 47. Methods that appear more than 5 times are shown in green, while those that appear fewer than 5 times are shown in red. If a method appears exactly five times, it is shown in white.
The interval forecast results at the significance level of $\alpha=0.05$ are available and can be viewed in a developed Shiny app at https://cfjimenezv.shinyapps.io/Shiny_App_CDF/.
6.6 Forecasting Age-Specific Survival Probabilities and Life Expectancy
The enhanced forecast accuracy of subnational life-table death counts ($\widehat{d}_{x}$) enables precise estimation of key demographic indicators, such as the age-specific probability of dying ($q_{x}$) and life expectancy ($e_{x}$). These metrics are derived by constructing a complete life table from the forecasted death counts ($d_{x}$). Specifically, we apply the standard life table relationships using the LifeTable and convertFx functions from the MortalityLaws package (Pascariu, 2025) in [Uncaptioned image]. The life table radix ($l_{0}$) is set to 100,000 for all forecasted tables. For illustration, Figure 8 shows the forecasted distributions of deaths, probabilities of dying, and life expectancy for the prefecture Hokkaido for the years 2007 to 2023, based on data from 1975 to 2006.
Figure 8: Forecasted life table metrics for Hokkaido, 2007–2023. Panels show age-specific deaths ($d_{x}$), probability of dying ($q_{x}$), and life expectancy ($e_{x}$) for females (F) and males (M).
6.7 Model Diagnostics
We illustrate model diagnostics using the FANOVA residuals to assess the weak functional white noise assumption, focusing on stationarity and conditional heteroscedasticity. Stationarity was tested using the $\mathbf{T}$-statistic-based method for functional time series (T_stationary in ftsa package in [Uncaptioned image], Hyndman & Shang, 2025), separately for female and male residuals (matrices $110\times 49$ per prefecture). Conditional heteroscedasticity was assessed using the functional conditional heteroscedasticity test (fCH_test in FTSgo package in [Uncaptioned image], Rice et al., 2020). Table 3 summarizes the $p$-values across the 47 prefectures. Since all $p$-values exceed 0.05, the residuals can be considered approximately stationary. Small $p$-values indicate some heteroscedasticity, but as our methods model the conditional mean, variance modeling is beyond the scope; functional GARCH-type models could be explored in future work.
Table 3: Summary of $p$-values from stationarity and conditional heteroscedasticity tests.
   Gender    Min    1st Qu.    Median    Mean    3rd Qu.    Max
Stationary test
   Female    0.060    0.078    0.085    0.0853    0.091    0.108
   Male    0.057    0.0705    0.079    0.0784    0.0855    0.099
Conditional heteroscedasticity test
   Female    0.0000    8.22e-15    1.64e-12    7.67e-04    2.58e-10    3.60e-02
   Male    3.85e-13    3.35e-09    6.56e-06    1.61e-02    2.13e-04    4.88e-01
7 Conclusion
We present several graphical tools, namely image plots and cross-correlation plots, to visualize the differences and dependence in life-table death counts between each prefecture and Japan as a whole.
Via the CDF transformation, we consider a suite of functional time-series models to model and forecast the subnational age distribution of death counts for 47 Japanese prefectures. The CDF transformation is ideal for handling zero values in subnational data. Within the CDF transformation, we also present two general strategies for constructing pointwise prediction intervals.
By dividing the datasets into training and testing samples, we evaluate and compare the point and interval forecast accuracy among the functional time-series models. We confirm the importance of joint modeling techniques, in which the MFTS, MLFTS, FANOVA, and HDFPCA models generally perform well on the aggregate level across 47 prefectures. We present a novel heatmap showing the frequency (of 47 prefectures), in which one method performs the best for each forecast horizon. Individual forecast errors for various horizons, obtained from all methods for each prefecture, are available in a developed shiny app at https://cfjimenezv.shinyapps.io/Shiny_App_CDF/. The forecast life-table death counts are not only important to compute age-specific survival probability and life expectancy in demography, but they are also useful to determine annuity prices for various ages and maturities studied in Shang & Haberman (2020).
There are several ways in which the methodology can be extended, and we briefly list five. 1) The functional standard deviation in (8) was computed coordinate-wise. There is a range of functional depth that can be implemented to compute other variants of standard deviations. 2) Instead of symmetric prediction intervals, asymmetric ones can be considered, searching for two tuning parameters to adjust the lower and upper bounds. 3) Third, the data set was divided almost equally into training, validation, and testing samples. Other proportions may be possible, leading to a more accurate selection of the tuning parameter $\xi_{\alpha}$ and more accurate interval forecasts. 4) We implemented a suite of functional time-series models for evaluation and comparison. Other time-series extrapolation models may also be considered, such as the robust functional principal component analysis by Oguamalam et al. (2025). It extends the Mahalanobis distance to Bayes spaces and introduces a regularized approach for improved principal component estimation in the presence of outliers. 5) We focus on modeling and forecasting the conditional mean. For modeling and forecasting conditional heteroskedasticity, the functional generalized autoregressive conditional heteroskedasticity (GARCH) model may be more adequate (see, e.g., Aue et al., 2017; Cerovecki et al., 2019; Rice et al., 2020 and 2023; Kearney et al., 2025). Future work will be to incorporate the GARCH framework into some functional models considered.
Supplementary Materials
  1. 1.
    Code for the forecasting methods based on the CDF transformation.The [Uncaptioned image] code to produce point and interval forecasts from the different functional time-series models described in the paper. The [Uncaptioned image] codes are available at the following repository: https://github.com/cfjimenezv07/HDFTS_CDF/tree/main/Rcode
  2. 2.
    Code for a shiny application.The [Uncaptioned image] code to produce a shiny user interface for the point and interval forecasts for the 47 Japanese prefectures. The [Uncaptioned image] codes are available at the following repository: https://github.com/cfjimenezv07/HDFTS_CDF/tree/main/Shiny_App_CDF
Acknowledgment
This research was supported by the Australian Research Council Discovery Project (grant no. DP230102250) and Future Fellowship (grant no. FT240100338). We are grateful for the many insightful comments and suggestions from the seminar participants at the University of Melbourne and the University of New South Wales. The author is grateful to Dr Yang Yang from the University of Newcastle for assistance with the heatmaps.
Appendix A: Point forecast accuracy when $K=6$
In Table LABEL:tab:appendix_1, averaged across 47 prefectures, we compare the point forecast accuracy among the functional time-series models, when the number of retained components is set to six as in Hyndman et al. (2013). Based on the summary statistics of the KLD and JSD, the MFTS provides the most accurate point forecasts for the female and male data.
Table 4: Averaged across 47 prefectures, we evaluate and compare the point forecast accuracy of the functional time-series models, measured by the KLD and JSD. The number of components is determined by setting $K=6$.
Female Male
Metric $h$ UFTS MFTS MLFTS FANOVA UFTS MFTS MLFTS FANOVA
KLD 1 0.004 0.004 0.004 0.006 0.003 0.003 0.003 0.004
2 0.006 0.008 0.006 0.007 0.004 0.004 0.004 0.005
3 0.008 0.009 0.008 0.008 0.005 0.005 0.005 0.005
4 0.010 0.011 0.009 0.009 0.006 0.006 0.006 0.006
5 0.012 0.013 0.010 0.011 0.006 0.007 0.006 0.007
6 0.014 0.014 0.011 0.013 0.007 0.007 0.007 0.007
7 0.017 0.016 0.013 0.015 0.008 0.008 0.008 0.008
8 0.021 0.018 0.016 0.018 0.009 0.009 0.009 0.009
9 0.026 0.021 0.019 0.022 0.010 0.010 0.010 0.011
10 0.032 0.025 0.023 0.027 0.011 0.012 0.011 0.011
11 0.039 0.029 0.027 0.033 0.012 0.013 0.013 0.013
12 0.047 0.035 0.034 0.041 0.012 0.015 0.014 0.014
13 0.058 0.042 0.040 0.048 0.013 0.017 0.015 0.015
14 0.067 0.038 0.047 0.063 0.014 0.011 0.015 0.018
15 0.088 0.053 0.067 0.083 0.014 0.014 0.020 0.023
16 0.116 0.071 0.089 0.110 0.015 0.018 0.027 0.030
17 0.151 0.085 0.106 0.125 0.015 0.020 0.031 0.033
Mean 0.042 0.029 0.031 0.038 0.010 0.011 0.012 0.013
Median 0.026 0.021 0.019 0.022 0.010 0.010 0.010 0.010
JSD 1 0.032 0.034 0.033 0.037 0.027 0.026 0.026 0.031
2 0.040 0.042 0.039 0.040 0.033 0.032 0.032 0.034
3 0.044 0.046 0.043 0.043 0.035 0.034 0.034 0.036
4 0.048 0.049 0.047 0.047 0.038 0.037 0.037 0.038
5 0.052 0.053 0.050 0.051 0.040 0.039 0.039 0.040
6 0.057 0.056 0.053 0.055 0.043 0.041 0.041 0.042
7 0.062 0.059 0.057 0.059 0.045 0.043 0.043 0.044
8 0.067 0.062 0.061 0.065 0.048 0.045 0.045 0.047
9 0.074 0.066 0.066 0.070 0.051 0.047 0.048 0.049
10 0.081 0.070 0.072 0.077 0.053 0.049 0.050 0.051
11 0.090 0.076 0.079 0.087 0.055 0.051 0.052 0.054
12 0.101 0.083 0.088 0.098 0.057 0.053 0.055 0.057
13 0.115 0.090 0.098 0.109 0.059 0.056 0.058 0.060
14 0.128 0.094 0.109 0.124 0.060 0.055 0.061 0.066
15 0.149 0.111 0.130 0.144 0.061 0.060 0.069 0.074
16 0.175 0.128 0.150 0.169 0.063 0.066 0.079 0.085
17 0.201 0.137 0.164 0.181 0.063 0.069 0.085 0.089
Mean 0.089 0.074 0.079 0.086 0.049 0.047 0.050 0.053
Median 0.074 0.066 0.066 0.070 0.051 0.047 0.048 0.049
Appendix B: Interval forecast accuracy when $K=6$
In Table LABEL:tab:appendix_2_K6, averaged across 47 prefectures, we compare the interval forecast accuracy among the functional time-series models. From the summary statistics of the CPD and interval scores at the level of significance $\alpha=0.2$, the MFTS offers the smallest CPD value for the female data, while the MLFTS offers the smallest CPD values for the male data. In terms of interval scores, the MFTS yields the smallest for the female data, while FANOVA yields the smallest for the male data.
Table 5: Averaged across 47 prefectures, we evaluate and compare the interval forecast accuracy, as measured by the CPD and interval score, among the functional time-series models at the level of significance $\alpha=0.2$. The number of components is determined by setting $K=6$. For either females or males, we highlight the method with the smallest overall CPD and interval scores.
Female Male
Metric $h$ UFTS MFTS MLFTS FANOVA UFTS MFTS MLFTS FANOVA
CPD 1 0.052 0.042 0.045 0.058 0.054 0.049 0.046 0.051
2 0.059 0.038 0.046 0.061 0.056 0.058 0.050 0.055
3 0.066 0.041 0.057 0.068 0.055 0.060 0.052 0.057
4 0.073 0.045 0.061 0.079 0.058 0.068 0.056 0.061
5 0.072 0.045 0.059 0.084 0.058 0.070 0.054 0.062
6 0.080 0.052 0.065 0.095 0.061 0.073 0.058 0.071
7 0.084 0.049 0.066 0.097 0.059 0.073 0.054 0.069
8 0.089 0.052 0.068 0.105 0.061 0.078 0.055 0.071
9 0.090 0.054 0.073 0.106 0.070 0.077 0.060 0.067
10 0.101 0.055 0.082 0.115 0.071 0.077 0.068 0.078
11 0.108 0.056 0.087 0.128 0.075 0.084 0.076 0.118
12 0.109 0.054 0.087 0.127 0.074 0.081 0.071 0.129
13 0.113 0.054 0.076 0.127 0.071 0.075 0.075 0.128
14 0.119 0.064 0.085 0.129 0.067 0.076 0.081 0.155
15 0.120 0.091 0.093 0.135 0.065 0.080 0.084 0.156
16 0.109 0.103 0.104 0.140 0.062 0.062 0.088 0.153
Mean 0.089 0.056 0.073 0.112 0.065 0.071 0.063 0.094
Median 0.090 0.053 0.067 0.108 0.062 0.073 0.058 0.071
$\overline{S}_{0.2}$ 1 233 253 234 248 239 221 208 225
2 281 292 275 301 268 252 237 251
3 326 325 310 352 294 278 261 275
4 374 357 346 406 324 305 285 299
5 421 388 376 460 348 327 304 321
6 485 433 421 533 381 354 330 350
7 558 480 468 614 403 372 353 369
8 649 531 533 721 428 392 380 395
9 734 587 602 825 450 408 404 418
10 845 659 689 961 472 436 429 438
11 969 756 800 1117 486 460 455 457
12 1134 848 896 1305 513 485 471 487
13 1283 872 999 1507 516 451 489 513
14 1508 1035 1220 1804 552 504 572 573
15 1834 1226 1466 2122 615 578 714 656
16 2130 1463 1754 2410 840 797 942 903
Mean 885 667 841 1376 484 465 502 460
Median 734 587 689 1305 472 436 455 457
Appendix C: Conformal prediction intervals
In machine learning, a popular methodology known as conformal prediction (Shafer & Vovk, 2008) is used to construct probabilistic forecasts calibrated on out-of-sample errors. From the absolute value of $\widehat{\varepsilon}_{1,s}^{g}(u),\dots,\widehat{\varepsilon}_{M,s}^{g}(u)]$, we compute its $100(1-\alpha)\%$ quantile for a level of significance $\alpha$, denoted by $q_{\alpha}(u)$. The prediction interval can be obtained as
$\left[\widehat{d}^{g}_{n+h|n,s}(u)-q_{\alpha}(u),\quad\widehat{d}^{g}_{n+h|n,s}(u)+q_{\alpha}(u)\right],$
where $\widehat{d}^{g}_{n+h|n,s}(u)$ denotes the $h$-step-ahead point forecasts for the data in the testing set.
The conformal prediction approach performs well when the validation set has a reasonably large sample size (Dhillon et al., 2024). For smaller sample sets, we recommend using the standard deviation method described in Section 5. As shown in Table LABEL:tab:appendix_3, the MLFTS achieves the lowest CPD values under the EVR criterion, whereas the MFTS provides the smallest CPD values for $K=6$ in the female dataset. For the male dataset, MLFTS is the preferred method. Regarding interval scores, MFTS attains the lowest values under the EVR criterion, while HDFPCA performs best for $K=6$ in the female data. For males, UFTS and FANOVA are favored under the EVR criterion, with MFTS performing best for $K=6$.
Table 6: Averaged across 47 prefectures, we evaluate and compare the interval forecast accuracy, measured by the CPD and interval score, for the functional time-series models at the significance level $\alpha=0.2$. The number of components is determined by the EVR criterion or $K=6$; for each, we highlight the method with the smallest overall CPD and interval score values.
EVR $K=6$
Metric Sex $h$ UFTS MFTS MLFTS FANOVA HDFPCA UFTS MFTS MLFTS FANOVA
CPD F 1 0.090 0.109 0.063 0.090 0.102 0.067 0.056 0.054 0.067
2 0.105 0.106 0.065 0.105 0.052 0.080 0.057 0.056 0.080
3 0.120 0.099 0.066 0.120 0.118 0.091 0.057 0.062 0.091
4 0.136 0.094 0.067 0.136 0.044 0.103 0.059 0.066 0.103
5 0.141 0.084 0.062 0.141 0.130 0.109 0.060 0.064 0.109
6 0.159 0.083 0.063 0.159 0.052 0.124 0.062 0.068 0.124
7 0.161 0.080 0.071 0.161 0.133 0.133 0.064 0.073 0.133
8 0.163 0.074 0.071 0.163 0.060 0.138 0.063 0.071 0.138
9 0.172 0.077 0.075 0.172 0.130 0.147 0.068 0.077 0.147
10 0.183 0.080 0.081 0.183 0.054 0.161 0.075 0.084 0.161
11 0.192 0.081 0.096 0.192 0.135 0.171 0.084 0.095 0.171
12 0.199 0.095 0.106 0.199 0.055 0.178 0.095 0.108 0.178
13 0.198 0.095 0.107 0.198 0.168 0.176 0.099 0.112 0.176
14 0.221 0.115 0.129 0.221 0.082 0.188 0.115 0.132 0.188
15 0.256 0.158 0.162 0.256 0.241 0.219 0.158 0.164 0.219
16 0.299 0.190 0.186 0.299 0.138 0.256 0.191 0.201 0.256
Mean 0.184 0.103 0.096 0.184 0.117 0.152 0.092 0.101 0.152
Median 0.177 0.094 0.078 0.177 0.118 0.136 0.071 0.081 0.136
M 1 0.094 0.122 0.062 0.094 0.090 0.058 0.064 0.063 0.058
2 0.113 0.127 0.072 0.113 0.111 0.070 0.077 0.068 0.070
3 0.119 0.130 0.076 0.119 0.115 0.076 0.083 0.070 0.076
4 0.125 0.137 0.083 0.125 0.117 0.083 0.091 0.077 0.083
5 0.123 0.137 0.087 0.123 0.122 0.090 0.098 0.080 0.090
6 0.133 0.146 0.097 0.133 0.137 0.105 0.109 0.090 0.105
7 0.134 0.139 0.099 0.134 0.142 0.111 0.107 0.090 0.111
8 0.126 0.132 0.095 0.126 0.139 0.117 0.112 0.089 0.117
9 0.116 0.127 0.093 0.116 0.139 0.118 0.110 0.087 0.118
10 0.119 0.133 0.096 0.119 0.154 0.119 0.117 0.092 0.119
11 0.114 0.129 0.100 0.114 0.160 0.118 0.118 0.099 0.118
12 0.110 0.130 0.102 0.110 0.177 0.122 0.118 0.103 0.122
13 0.100 0.121 0.093 0.100 0.172 0.121 0.114 0.091 0.121
14 0.104 0.130 0.105 0.104 0.187 0.125 0.125 0.099 0.125
15 0.115 0.143 0.118 0.115 0.196 0.133 0.136 0.114 0.133
16 0.134 0.183 0.151 0.134 0.234 0.156 0.179 0.151 0.156
Mean 0.120 0.135 0.097 0.120 0.142 0.104 0.107 0.092 0.104
Median 0.119 0.133 0.096 0.119 0.138 0.118 0.113 0.091 0.118
$\overline{S}_{0.2}$ F 1 286 423 247 286 331 253 256 237 253
2 338 459 285 338 323 308 297 280 308
3 393 489 319 393 411 360 331 316 360
4 457 521 354 457 349 419 366 353 419
5 514 548 386 514 470 474 398 386 474
6 593 573 429 593 393 551 438 436 551
7 671 601 477 671 549 632 481 486 632
8 784 627 536 784 451 741 521 549 741
9 897 660 600 897 624 855 570 617 855
10 1050 715 676 1050 477 1001 633 707 1001
11 1217 778 766 1217 735 1170 710 804 1170
12 1396 877 862 1396 513 1356 802 908 1356
13 1588 865 954 1588 968 1563 820 1003 1563
14 1872 1015 1181 1872 570 1843 994 1237 1843
15 2185 1199 1387 2185 1368 2153 1201 1469 2153
16 2528 1343 1552 2528 619 2475 1358 1665 2475
Mean 1092 731 740 1092 603 1095 630 742 1095
Median 840 644 638 840 495 928 546 677 928
M 1 259 280 221 259 251 224 222 209 224
2 281 309 247 281 286 249 253 238 249
3 296 335 268 296 314 272 278 262 272
4 312 359 287 312 334 291 303 284 291
5 320 378 301 320 353 311 325 302 311
6 337 399 319 337 383 337 350 325 337
7 351 412 337 351 413 361 367 347 361
8 366 424 359 366 438 385 383 372 385
9 379 435 378 379 462 404 396 390 404
10 394 457 397 394 504 420 420 412 420
11 414 475 426 414 548 440 444 439 440
12 440 507 445 440 626 468 472 456 468
13 456 454 454 456 699 485 432 464 485
14 500 498 545 500 839 528 481 551 528
15 562 555 661 562 977 586 543 657 586
16 609 600 744 609 1094 640 595 730 640
Mean 393 419 396 393 516 395 387 414 395
Median 372 418 378 372 450 400 375 381 400
Appendix D: Model Confidence Set (MCS) Analysis
We performed a rigorous statistical comparison of the point forecast accuracy across all competing models using the Model Confidence Set (MCS) procedure (Hansen et al., 2011). The MCS procedure identifies a set of superior models, $\mathcal{M}^{*}$, where the null hypothesis of Equal Predictive Ability (EPA) cannot be rejected at a given significance level $\alpha$.
The analysis was based on the point forecast errors measured by the Kullback-Leibler Divergence (KLD) for nine Functional Time-Series (FTS) models over 17 forecast horizons ($h=1$ to $h=17$). The procedure was applied to the forecast errors for 47 Japanese prefectures separately for male and female mortality data. We used a significance level of $\alpha=0.10$, a block-bootstrap approach with $B=2000$ resamples, and the $T_{\max}$ statistic.
The heatmaps in Figure 9 visualize the results of the MCS procedure. In each cell, the count represents the number of prefectures (out of 47) for which the corresponding model was included in the statistically superior set $\mathcal{M}^{*}$ at that horizon, demonstrating the robust predictive ability of the models across the regional series. For the female data, the MLFTS method consistently achieved statistical superiority across most prefectures at shorter horizons. Conversely, the HDFPCA method proved most robust at longer horizons. For the male dataset, the MFTS method was consistently included in the superior set $\mathcal{M}^{*}$ for both short and long forecast horizons.
(a) Female data
(b) Male data
Figure 9: Heatmaps summarizing the results of the Model Confidence Set (MCS) procedure. The analysis compares the point forecast accuracy, measured by the Kullback-Leibler Divergence (KLD), across nine functional time-series models over 17 forecast horizons ($h=1$ to $h=17$) for both female and male mortality data. Each cell value indicates the number of the 47 prefectures for which the corresponding model was included in the superior set, $\mathcal{M}^{*}$, at that specific horizon (each row sums to 47). Cells are color-coded based on the count: Green highlights models included in the superior set for six or more prefectures ($\geq 6$), Red for four or fewer prefectures ($\leq 4$), and White for exactly five.